### v1

# Do Students Write Better Post-AI Support?

Abstract
[Abstract] Academic writing increasingly involves multimodal tasks requiring students to integrate visual information and textual arguments. While generative AI (GenAI) tools, like ChatGPT, offer new pathways for supporting academic writing, little is known about how students’ GenAI literacy influences their independent multimodal writing skills or how chatbot interaction strategies (passive reactive vs. proactive scaffolding) impact learning. This study examined 79 higher education students’ multimodal academic writing performance using a comparative research design. Students completed writing tasks integrating visual data under two chatbot-assisted conditions (passive vs. proactive) and subsequently without AI assistance. Their writing performance was rigorously evaluated across five dimensions, including insightfulness, visual data integration, organisation, linguistic quality, and critical thinking. Ordinal logistic regression and correlation analyses revealed that higher levels of GenAI literacy significantly predicted stronger independent multimodal writing performance immediately after AI assistance removal, particularly for students using passive chatbots requiring active prompting. These results highlight the critical role of GenAI literacy and specific chatbot interaction strategies in shaping students’ capacities for independent multimodal academic writing. Our findings emphasise the need for purposeful integration of GenAI literacy training into curricula and balancing external scaffolding support with autonomous learning opportunities. This research offers valuable recommendations for educators leveraging AI-enhanced pedagogies to optimise student writing outcomes and technological engagement strategies.
keywords:
generative AI literacy, academic writing, chatbot interaction, multimodal, skill development*Practitioner Notes
What is currently known about this topic
Generative AI (GenAI) tools are increasingly used by learners in their academic writing tasks, offering new opportunities and challenges.
Effective use of advanced AI tools depends partly on students’ technological literacy and their ability to strategically engage with AI-generated content.
The academic writing process is becoming increasingly multimodal, requiring integration of textual and visual data, yet limited research exists on how GenAI tools influence learner performance in such multimodal contexts.
What this paper adds
Provides empirical evidence that learners’ GenAI literacy significantly predicts their independent multimodal academic writing performance, especially after AI supports are removed.
Demonstrates that passive chatbot interactions benefit students with higher GenAI literacy more strongly than proactive chatbot interactions, highlighting interaction style as a crucial moderating variable.
Confirms GenAI literacy as a distinct predictor of writing performance, separate from domain-specific knowledge, indicating its importance for successfully integrating visual information and critical-thinking components into AI-assisted writing tasks.
Implications for practitioners
Educators should explicitly integrate GenAI literacy training within academic curricula, focusing on skills like formulating effective prompts, critically evaluating AI-generated output, and independently integrating this assistance into academic work.
Practitioners need to carefully select chatbot scaffolding strategies appropriate for students’ GenAI literacy, balancing proactive guidance with opportunities for learners’ active inquiry, metacognitive awareness, and autonomous problem-solving.
Instructional designers should strategically balance GenAI-driven scaffolding with independent practice opportunities for students to independently apply and develop their GenAI literacy skills.
1 Introduction
Academic writing is an essential practice in higher education. It helps students develop critical thinking, effective communication, and the ability to deal with complex ideas and information 33, 46. Students need to interpret and synthesise data, create logical arguments, and assess different perspectives 13, 47. Increasingly, academic writing also involves multimodal literacy, requiring students to accurately interpret, synthesise, and incorporate varied visual data sources, such as graphs, charts, or images, into their arguments. This skill reflects the increasingly multimodal nature of contemporary academic discourse and demands additional capabilities concerning visual-data integration 34, 46, 27. Academic writing is crucial for both academic success and professional competency 53, 59. Employers value strong academic writing skills as they help individuals write persuasive reports, create coherent presentations, and solve problems effectively 12. Yet, academic writing is linked to cognitive and metacognitive processes. It encourages students to monitor and control their thought processes, facilitating metacognitive awareness and self-regulation 18. This engagement is essential for students’ ability to plan, evaluate, and reflect on their learning strategies, leading to a deeper understanding and better retention of information 21.
As technology advances, academic writing can evolve from being solely an individual process to one supported by various tools 45. With the increasing availability of generative AI (GenAI) tools (e.g., ChatGPT and Gemini), students have more opportunities to enhance their academic writing performance 62, 4. Using GenAI tools in writing is becoming more common 30, making it urgent to understand how these external supports impact students’ skill development 39. GenAI tools can provide feedback, generate content ideas, or assist in structuring arguments, offering new ways for students to engage with writing tasks 7. However, while these tools present exciting possibilities, it is critical to examine their influence on learners’ development of independent writing skills 63. Understanding this dynamic will guide educators in integrating technology effectively into writing instruction, ensuring that while students benefit from technological aids, they also continue to grow their critical thinking and independent writing abilities 45, 7.
Indeed, emerging studies are beginning to explore how GenAI can produce high-quality writing and support student writing processes 38, 30, 39. Most of these studies focus on how GenAI can assist humans in writing and the patterns that emerge from this interaction 45, 62. However, existing research predominantly addresses text-only writing contexts, leaving uncertainties about the role of GenAI in supporting multimodal academic writing involving visual information 27, 34. There is also little empirical evidence on how GenAI tools affect students’ academic writing performances, especially when these tools are taken away after they have been used. Using GenAI tools might merely create the illusion of performance enhancement, but it might not necessarily lead to learning and skill improvements 16, 63. Additionally, few studies have thoroughly examined how different interaction strategies of GenAI, such as passively replying versus proactively scaffolding, impact student writing. Understanding interaction strategy is important because it could shape how students engage with these tools and influence the development of their writing skills 45, 29. A passive interaction might mean the tool only responds to commands, while proactively scaffolding could guide students step-by-step and provide feedback 63. Each style can affect how independent and critical students become in their writing. Thus, it is crucial to understand if dependence on GenAI affects students’ writing skills and independence. Research in this area can help educators balance the use of GenAI technology, making sure students can take advantage of GenAI while still growing their own academic writing abilities 45, 16, 7.
The current study contributes empirical evidence to address the aforementioned gaps. Specifically, we conducted a comparative study with 79 higher education students who completed an academic writing task based on visual data, with the support of a passive (reactive replying) or proactive (active scaffolding) GenAI agent. Additionally, students completed a similar multimodal academic-writing task with visual information integration but with GenAI support removed, to examine potential lasting effects. Students’ written responses were rigorously scored using a comprehensive rubric covering insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking. Furthermore, we investigated how students’ generative AI literacy influenced their multimodal academic writing performance when provided with and subsequent to the removal of GenAI assistance. By addressing these points, our study contributes novel insights into the differential impacts of passive versus proactive interaction strategies on students’ abilities to integrate textual and visual data, and underscores the role of learners’ varying GenAI literacy in determining academic writing performances. These findings extend existing research by offering practical implications for educators and practitioners aiming to integrate GenAI effectively into multimodal academic writing practices.
2 Background
2.1 Academic Writing
Academic writing is a core skill in higher education, requiring students to effectively synthesise ideas, engage critically with evidence, and communicate coherently. At the heart of successful academic writing lies the ability to provide insightful interpretations by deeply analysing the subject matter, demonstrating originality, and making sophisticated connections among multiple sources 32, 59. Beyond text alone, contemporary academic discourse increasingly involves multimodal literacy, in particular through the integration of visual data. Students must accurately interpret, meaningfully synthesise, and seamlessly incorporate varied graphical sources into their narrative 34, 46, 27. Equally central is the structural organisation and coherence of written arguments, allowing ideas to flow logically and enhancing reader comprehension 23, 53. Further, linguistic quality, marked by appropriate grammar, syntax, vocabulary, and academic style, is indispensable, ensuring precision, clarity and credibility within scholarly communication 5, 24. Lastly, academic writing requires rigorous critical thinking, encompassing analytical evaluation, reasoned argumentation, and thorough exploration of competing ideas or solutions 43, 15.
Assessing students’ academic writing holistically thus demands evaluation across these multiple dimensions. Insightfulness is typically measured through evidence of original, nuanced interpretations rather than superficial summaries 32, 9. Visual data integration is judged by examining the accuracy and relevance of references to graphical material, as well as how effectively visual information enhances overall arguments 27. Organisation and coherence involve assessing logical structuring and the clarity with which ideas connect and flow progressively 47. Linguistic quality is commonly evaluated through accuracy in grammar, usage, and stylistic appropriateness, reflecting scholarly conventions and readability 5. Finally, students’ critical thinking skills are commonly demonstrated through their ability to carefully analyse evidence, balance perspectives, and propose reasoned solutions or arguments based on comprehensive evaluation 15, 43. Combined, these dimensions offer a nuanced, comprehensive assessment of students’ academic writing performance, which is important for capturing students’ academic success and intellectual growth.
2.2 Generative AI and Writing
[…]
2.3 Generative AI Literacy
Given the increasing use of GenAI in educational contexts, particularly in academic writing, it has become essential to investigate and consider students’ GenAI literacy when exploring AI-assisted writing practices. While general AI literacy has been defined and explored extensively in previous frameworks 40, 44, the unique characteristics of generative technologies necessitate a more specialised literacy framework. Researchers have highlighted that GenAI literacy involves not only understanding how generative tools function, but also mastering practical skills to effectively prompt these tools, critically evaluate AI-generated content, and navigate ethical and contextual complexities specific to GenAI-generated outputs 6, 69. Despite recognition of the importance of GenAI literacy for meaningful interaction with tools like ChatGPT and Gemini, there remains limited empirical research that explicitly examines how students’ varying levels of GenAI literacy impact their academic writing performances when GenAI assistance is provided or withdrawn 3, 41. Addressing this gap is imperative, as understanding the relationship between students’ GenAI literacy and their academic writing outcomes can inform practical strategies for educators seeking to effectively integrate these advanced technologies into multimodal academic writing instruction 11, 3.
2.4 Domain Knowledge
Domain knowledge is an essential consideration when investigating AI-assisted multimodal academic writing, particularly when visual information and interpretation are integral to the task. Students’ capacity to successfully integrate visual data into meaningful written discussions depends heavily on their familiarity with relevant subject areas 1, 55. This underscores the necessity of evaluating and accounting for the domain knowledge possessed by individual learners, especially when examining the relationship between their GenAI literacy and writing performance 2. Without considering domain proficiency, researchers may inadvertently attribute students’ challenges in producing coherent multimodal arguments or effectively utilising GenAI supports to deficiencies in GenAI literacy alone, rather than identifying underlying gaps in subject-specific knowledge 1. Accordingly, research into GenAI literacy for multimodal academic writing contexts should adopt a nuanced approach, explicitly acknowledging the role of domain-specific expertise alongside GenAI literacy.
2.5 Research Gaps
The current study aimed to address the aforementioned gaps by examining how students’ GenAI literacy affects their multimodal academic writing performance with and without GenAI support. Specifically, we investigated whether these effects differ based on chatbot interaction strategies (passive vs. proactive) and account for students’ domain-specific knowledge to ensure an accurate interpretation of the effects of GenAI literacy. Specifically, the following research questions were investigated:
RQ1: What is the relationship between students’ GenAI literacy and their academic writing performance at the AI-support and AI-removal stages, and does this relationship differ based on chatbot interaction type (passive vs. proactive)?
RQ2: Do students with low versus high GenAI literacy differ significantly in their academic writing performance at the AI-support and AI-removal stages, and do these differences differ based on chatbot interaction type (passive vs. proactive)?
RQ3: To what extent does students’ GenAI literacy predict their academic writing performance at the AI-removal stage after controlling for domain-specific knowledge, chatbot interaction type (passive vs. proactive), and writing performance at the AI-support stage?
3 Methods
3.1 Learning Contexts
The learning context involved writing based on visual information, a critical skill in academic writing, closely aligned with tasks such as IELTS Academic Writing Task 1 25. Participants received three distinct visual representations (a bar chart, a network diagram, and a ward map; described in detail in Section 3.3.1). They were instructed to interpret these visuals and compose a short written answer (150 words) addressing a question about team performance (e.g., "How the two nurses managed the primary patient (Amy) while attending to other beds, focusing on their task prioritisation, verbal communication, and stress levels?"). The visuals were specifically designed using data from authentic learning contexts, such as healthcare simulations, ensuring participants had no prior exposure to these materials. Thus, any potential effects stemming from participants’ prior familiarity with the presented information were minimal. The study was conducted through an online environment (details in Section 3.3.2).
3.2 Participants
A total of 81 participants completed the study. However, two participants were excluded from the analysis due to providing written responses shorter than 20 words, resulting in a final sample of 79 participants. This exceeded the minimum sample size (72 participants; 36 per condition) required by a priori power analysis (effect size , power = 0.80, ). Participants were recruited via Prolific111https://www.prolific.co/ and compensated £8 for approximately one hour of participation. Participants were randomly assigned to one of two experimental conditions: passive chatbot (39) or proactive chatbot (40). All participants were current students or graduates from medical or nursing programs. The sample comprised 45 females and 34 males, primarily aged 18-24 (28) and 25-34 (31); fewer participants belonged to age groups 35-44 (11), 45-54 (6), 55-64 (2), and 65 or older (1). Participants were distributed geographically across North/Central America (34), Europe (21), Africa (15), Australia (4), South America (2), and other regions (3). Educational backgrounds included Bachelor’s degrees (36), high school diploma or equivalent (18), Master’s degrees (11), vocational training or diplomas (7), Doctoral degrees (4), and other qualifications (3). Familiarity with GenAI tools was predominantly intermediate (42), beginner (25), advanced (9), expert (2), and none (1).
3.3 Materials and Apparatus
3.3.1 Learning Materials
The three visual information representations used for learning were generated from data collected during a high-fidelity healthcare simulation. In this simulation, students worked in teams of four (comprising two primary and two secondary nurses) to navigate acute clinical scenarios within a realistic, ward-like environment. The simulated setting included hospital beds, medical equipment, and manikins designed to replicate patient vital signs. The objective of the simulation was to enhance students’ prioritisation skills, communication strategies, and collaborative teamwork by replicating realistic clinical emergencies. Throughout the simulation, students’ positional information (x-y coordinates and body orientation) was captured using an indoor positioning system; audio interactions were recorded through wireless headset microphones; and physiological data, specifically heart rates, were collected using Fitbit Sense wristbands.
The multimodal data were used to generate three visuals (Figure 1), each with progressively increasing complexity. First, a bar chart summarises time spent on different behaviours derived from positional data. This visual information illustrates team-level resource allocation and prioritisation strategies 64. Second, a communication network diagram combining positional and audio data to reveal the frequency and direction of verbal interactions among nurses, the doctor, the patient, and the patient’s relative, thereby depicting participant roles and their level of involvement 68. Third, a ward map integrates positional, audio, and physiological data. Inspired by visual and spatial analytics methods used in professional sports 20, the ward map utilises heatmap visualisation techniques to show communication intensity, trace spatial movements, and highlight locations of peak physiological arousal via heart rate data.
3.3.2 Research Platform
The academic writing tasks were conducted using a custom-built platform comprising three main components. The first component, the display area (Figure 2), presented one visual analytic at a time to minimise visual clutter and prevent information overload 14. Participants could freely navigate between the three visual information representations using the navigation arrows positioned on either side of the component. The second component, the writing space (Figure 2), provided participants with an area to articulate their analytical insights by synthesising information presented in the visual information. The third component varied based on the experimental condition. In the AI-supported conditions, participants had access to a chat function (Figure 2), offering real-time support from GenAI agents, which responded either passively or proactively. Conversely, in conditions without AI support, the chat function was replaced by an instruction component providing standard task instruction without AI assistance. The overall platform design emphasised simplicity 52, enabling participants to focus primarily on comprehending the visual information and effectively completing the tasks without being overwhelmed by complex functionalities.
3.3.3 Chatbot Interaction Design
[…]
Building upon the VizChat prototype 65, both chatbots adopted an advanced agentic architecture composed of four interconnected modules. Prompts were categorised into passive and proactive types corresponding to the aforementioned interaction strategies. At its core, the GenAI agent managed and synthesised inputs from visual information and user prompts, utilising Retrieval-Augmented Generation (RAG) to efficiently retrieve data-informed content from its knowledge database 19. This comprehensive knowledge repository contained contextual information, visualisation details, and interaction logs that enabled personalised and relevant interactions. Finally, synthesised prompts were processed by the multimodal GPT-4o AI model, generating coherent and context-aware responses that enriched the database and progressively improved the quality of subsequent learner interactions.
3.4 Measures and Instruments
3.4.1 Generative AI Measurement (GLAT)
We assessed learners’ generative AI literacy using the Generative AI Literacy Assessment Test (GLAT) 28. This 20-item instrument was selected for its established validity and reliability. Face validity was ensured through expert review and revision, while content validity was evaluated with 204 participants to confirm item relevance, comprehensiveness and comprehensibility. Structural validity and internal consistency were confirmed using classical test theory and item response theory, supporting a reliable two-parameter logistic (2PL) model (Cronbach’s ) and robust factor structure (, , , 90% CI ). As the 2PL model accounts only for item difficulty and discrimination, we applied a correction-for-guessing formula 54 to control for potential guessing effects (Eq. 1).
In this formula, represents the final score corrected for guessing behaviours, denotes the number of correctly answered items by a given participant, denotes the number of wrongly answered items, and denotes the number of answer choices per item, which is four for the GLAT.
3.4.2 Domain Knowledge Measurement (VLAT)
Visualisation literacy constitutes essential domain knowledge for effective multimodal writing 17, 37. We assessed participants’ domain knowledge in visualisation literacy using the Visualisation Literacy Assessment Test (mini-VLAT) 48. This 12-item instrument was selected for its established validity and reliability, which are comparable to those of the original 53-item VLAT 37, while providing a briefer assessment suitable for our extended study. Its concise structure helped minimise participants’ cognitive load. Structural validity and internal consistency for this instrument have previously been confirmed 48. To account for potential guessing effects, we also applied the correction-for-guessing formula 54 to participants’ scores.
3.4.3 Academic Writing Assessment Rubric
A comprehensive assessment rubric was developed to evaluate the participants’ academic writing performances in tasks involving visual data analyses. Based on prior literature and established theoretical grounding within academic writing research (as elaborated in Section 2.1), we identified five distinct evaluation metrics (Table 1): Insightfulness,Visual Data Integration, Organisation and Coherence, Linguistic Quality, and Critical Thinking. Each criterion was evaluated using a 5-point Likert scale (Level 1–Poor to Level 5–Excellent), with comprehensive descriptors and examples available in the full rubric (see Supplementary Information).
| Metric | Description | Reliability () |
| Insightfulness | Assesses the originality, depth, and relevance of insights. High-quality responses identify subtle patterns, offer holistic interpretations, and provide conclusions that extend beyond obvious points. Low-quality responses are surface-level observations or simplistic interpretations 32, 9. | 0.82 |
| Visual Data Integration | Measures the effectiveness and clarity with which visual data (charts, graphs, communication networks, heart rate information) is explicitly interpreted, combined, and connected into the written analysis. High-quality responses synthesize multiple visual data sources cohesively, while low-quality responses merely mention data without synthesizing or miss integrating data entirely 27. | 0.85 |
| Organisation and Coherence | Evaluates the clarity, logical structure, and connectedness of ideas presented in the writing. High-quality responses have logical structure, smooth transitions between points, and clearly organized arguments. Low-quality responses lack clarity, exhibit abrupt transitions, or are disorganized 47. | 0.75 |
| Linguistic Quality | Reflects grammar accuracy, vocabulary appropriateness, sentence fluency, and overall clarity of written expression. High-quality responses utilize clear, precise, and varied language with minimal errors, while low-quality responses contain grammatical errors, limited vocabulary, or unclear expression 5. | 0.89 |
| Critical Thinking | Evaluates the level of careful thought, complexity, consideration of alternative viewpoints, and strength of conclusions and recommendations. High-quality responses deeply analyze complexities and implications, considering broader contexts, contingencies, or solutions. Low-quality responses show limited or superficial analytical engagement and fail to address alternative perspectives or implications 15, 43. | 0.80 |
3.4.4 Scoring Procedures and Inter-Rater Reliability
The scoring process followed detailed guidelines based on the rubric provided in the Supplementary Information. Two independent, trained coders rated each participant’s writing samples at each study phase (AI-support and AI-removal). Coders were trained extensively on the rubric definitions through practice scoring sessions with sample responses (n=10), followed by discussion and calibration of understanding to maximise rating consistency. All essays were then independently blind-rated by both coders. After ratings were fully completed, reliability analyses were performed (Spearman’s correlations), demonstrating high inter-rater reliability ( in all cases; Table 1). Final participant scores for each dimension, at each stage, were averaged across the two coders, producing a single composite rating for each dimension and subsequently an overall Composite Score. To ensure transparency and replicability, example excerpts representing each scoring level were provided in the detailed rubric (see Supplementary Information).
3.5 Study Procedure
We conducted a two-phase, mixed-method experimental study to examine how GenAI literacy and AI interaction type (passive vs. proactive chatbot) influenced participants’ academic writing performance with visual information. Participants were recruited via Prolific and directed to a Qualtrics222https://www.qualtrics.com/ survey, where they provided informed consent, demographics, and completed literacy assessments (GLAT and mini-VLAT). After reviewing contextual information on the healthcare simulation (Section 3.3.1), they were randomly assigned to either the passive or proactive chatbot assistance condition. In the AI-supported phase, participants composed a 150-word written response interpreting the provided multimodal visuals with real-time chatbot support. Subsequently, in the AI-removal phase, participants completed a similar task with new visual data from a different simulation scenario without chatbot access (Figure 4). The study lasted approximately one hour. Ethical approval was obtained from Monash University (Project Number: 37307).
3.6 Analysis
3.6.1 Correlation Analysis (RQ1)
To examine the relationship between participants’ GenAI literacy and academic writing performance (RQ1), correlation analyses were conducted separately for each chatbot engagement condition (passive, proactive), at both AI-support and AI-removal time points, and across the five writing dimensions, as well as the composite scores. GenAI literacy was first screened for normality to determine the suitable statistical test. The Shapiro-Wilk normality tests indicated a violation of normality assumptions for GenAI literacy (). Thus, Spearman’s rank-order correlation was used for all analyses. For each condition separately, Spearman’s correlations were calculated between participants’ GenAI literacy scores and their average writing scores at the AI-support and AI-removal stages.
3.6.2 Comparative Analysis (RQ2)
To investigate further if students at varying levels of GenAI literacy performed differently in academic writing (RQ2), we conducted comparative analyses by grouping participants into two distinct GenAI literacy categories: low and high literacy. Participants were categorised based on median splits of GenAI literacy scores within each condition separately. We chose the median split procedure to ensure balanced group sizes and improve interoperability. To ascertain whether academic writing performance differed significantly between low and high GenAI literacy groups, we used Mann-Whitney U tests due to the ordinal nature of the rating scales and violations of normality assumptions (verified through Shapiro-Wilk tests). Analyses were conducted separately for each evaluation criterion (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) as well as for the composite score, for both the AI-support and AI-removal stages periods. Statistical significance was set at and adjusted for multiple comparisons using the Bonferroni method. Significant differences were further interpreted using effect sizes calculated as the Rank-Biserial correlation.
3.6.3 Regression Analysis (RQ3)
To investigate the extent GenAI literacy predicted writing performance after controlling for domain-specific knowledge (RQ3), ordinal logistic regression analyses were conducted. As shown in the following formula (1), the dependent variables for regression analyses were participants’ academic writing performance scores during the AI-removal period for each dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) separately, as well as the composite writing score. Each outcome variable, measured using ordinal scales (; scores from 1 to 5; refers to each ordinal threshold, ), justified the use of ordinal logistic regression rather than linear regression. Primary predictor variables included standardised GenAI literacy scores (), domain knowledge (; VLAT score), chatbot intervention type (; binary-coded: 0 = passive chatbot, 1 = proactive chatbot), and participants’ AI-assisted scores (; to control writing status during chatbot support availability). Additionally, denotes the threshold/intercept parameter for each category boundary, and , and represent regression coefficients for the predictors.
Prior to analysis, GenAI literacy scores were standardised (M = 0, SD = 1) to facilitate regression coefficient interpretation. Ordinal logistic models were fitted using a logit link function and estimated with Maximum Likelihood Estimation via the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm in the Python Statsmodels library. Before analyses, multicollinearity was examined through variance inflation factor (VIF) values, and proportional odds assumptions were evaluated using the Brant test. Statistical significance of model predictors was evaluated at the level. Regression results were interpreted by examining regression coefficients (), standard errors (SE), test statistics (z-values), significance levels (-values), and odds ratios (ORs). Odds ratios greater than one indicated an increased likelihood of scoring higher on the writing assessment due to higher levels of GenAI literacy, whereas odds ratios less than one indicated a decreased likelihood.
4 Results
4.1 Correlation Between GenAI Literacy and Writing Performance (RQ1)
Spearman’s rank-order correlations were conducted separately by chatbot engagement condition (Passive: ; Proactive: ) to assess the relationship between GenAI literacy and writing rubric scores at both AI-support and AI-removal stages (Figure 5). The analysis revealed significant positive correlations in the passive chatbot condition (Passive group) at the AI-removal stage, indicating that higher GenAI literacy scores were associated with higher academic writing performance after removing AI support. Specifically, GenAI literacy significantly correlated with Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and Composite Score (). Additionally, a significant positive correlation was found for the Passive group at the AI-support stage only for Linguistic Quality (). No statistically significant correlations between GenAI literacy and writing rubric metrics were observed within the proactive chatbot condition.
4.2 Comparisons Between Low and High GenAI Literacy Groups (RQ2)
To further explore differences in writing performance based on GenAI literacy levels, participants in the passive chatbot condition () were grouped into low () and high () literacy categories based on median splits of GenAI literacy scores. Mann-Whitney U tests were performed separately for each rubric dimension at both the AI-support and AI-removal stages. Medium to large differences emerged primarily at the AI-removal stage (without AI support), with high GenAI literacy participants consistently scoring higher across several writing dimensions. Specifically, at the AI-removal stage, significant differences were observed for Insightfulness (), with the high literacy group achieving higher scores (Mdn = 2.75, IQR = 1.00) compared to the low literacy group (Mdn = 2.00, IQR = 0.00); Visual Data Integration (), with higher scores for the high literacy group (Mdn = 2.50, IQR = 0.50) versus low literacy group (Mdn = 2.00, IQR = 0.00); and Organisation and Coherence (), again favouring high literacy (Mdn = 2.75, IQR = 0.50) over low literacy participants (Mdn = 2.00, IQR = 0.50). Additionally, a significant difference was found for the overall Composite Score (), with higher scores among high literacy (Mdn = 13.25, IQR = 3.38) than low literacy participants (Mdn = 11.00, IQR = 1.50). At the AI-support stage (with GenAI assistance), only Linguistic Quality revealed a significant difference (), with the high literacy group scoring higher (Mdn = 3.00, IQR = 0.50) compared to the low literacy group (Mdn = 3.00, IQR = 0.00). On the other hand, participants in the proactive chatbot condition () were similarly grouped into low () and high () GenAI literacy categories based on median splits of GenAI literacy scores. However, Mann-Whitney U tests revealed no significant differences between the low and high literacy groups for any of the rubric metrics at either the AI-support or AI-removal stages.
4.3 Ordinal Logistic Regression Predicting Writing Performance (RQ3)
For each ordinal logistic regression model, proportional odds assumptions were met (Brant test, all ), and multicollinearity diagnostics indicated no problematic multicollinearity (variance inflation factor [VIF] < 5 for all predictors). Results demonstrated that higher GenAI literacy significantly predicted superior writing performance after AI assistance was withdrawn for three rubric criteria: Visual Data Integration, , , , , ; Critical Thinking, , , , , ; and overall writing quality measured through the Composite Score, , , , , . Specifically, each one-unit increase in standardised GenAI literacy was associated with a 14% greater likelihood of achieving higher ratings in Visual Data Integration, a 15% greater likelihood for Critical Thinking, and an 11% greater likelihood to receive higher overall Composite Score. However, GenAI literacy did not significantly predict scores on the Insightfulness, Organisation and Coherence, and Linguistic Quality dimensions. Additionally, participants’ AI-assisted writing scores significantly predicted their subsequent writing performance across all rubric dimensions: Insightfulness, , , , ; Visual Data Integration, , , , ; Organisation and Coherence, , , , ; Linguistic Quality, , , , ; Critical Thinking, , , , ; and overall Composite Score, , , , . Neither domain-specific knowledge (VLAT scores) nor chatbot interaction type (proactive vs. passive) significantly predicted writing performance in any regression models. Thus, the predictive relationship of GenAI literacy with writing performance was robust regardless of participants’ domain-specific knowledge or chatbot intervention type.
5 Discussion
This study explored the association between GenAI literacy and students’ multimodal academic writing performance in conditions where either passive or proactive GenAI tools supported academic writing. By examining students’ independent writing performance immediately following the withdrawal of GenAI assistance, we sought to understand whether and how GenAI literacy shapes writing performance once external scaffolding is removed and how interactions with chatbots of different styles might moderate these relationships.
5.1 GenAI Literacy and Independent Writing Performance (RQ1)
For RQ1, our findings indicated a significant positive relationship between students’ GenAI literacy and their academic writing performance at the independent writing stage immediately after GenAI assistance was removed. Specifically, participants with higher GenAI literacy demonstrated stronger performance across multiple rubric dimensions (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking, and composite scores) when writing without AI support. This result provides empirical evidence underscoring the role of GenAI literacy as potentially influential in learners’ capabilities, particularly when performing independently without technological scaffolds. It appears that GenAI literacy is associated with superior performance in conditions where no external AI prompts or support mechanisms are immediately available. This finding is in line with prior theoretical positions suggesting that effective interactions with technology depend not only on the availability of tools but also on users’ understanding of how best to integrate and leverage them effectively 45, 63. Researchers argue that higher levels of literacy related to emerging digital tools lead to better competence, critical engagement with technology-generated output, and stronger independent strategic thinking 28, 30. Our results reinforce this perspective in the context of GenAI-assisted multimodal writing, highlighting that students with higher GenAI literacy appear better prepared, at least immediately following GenAI assistance, to produce high-quality, independent academic writing without continual external support.
5.2 The Role of Chatbot Interaction Strategies (RQ2)
Interestingly, for RQ2, significant relationships between higher GenAI literacy and independent academic writing performance emerged only within the passive (reactive) chatbot condition. In contrast, the proactive (actively scaffolding) chatbot condition did not demonstrate a significant correlation between GenAI literacy and subsequent independent writing performance. This finding suggests that chatbot interaction strategies can potentially moderate the relationship between GenAI literacy and student outcomes. One plausible explanation is that passive chatbots, which respond only when students actively query them, require greater cognitive and strategic engagement from learners. GenAI-literate students in such conditions may have had opportunities and abilities to formulate effective prompts, critically evaluate the AI-generated assistance, and transfer the insights gained into their independent writing tasks. Conversely, proactively scaffolding chatbots continuously direct and structure learner engagement, perhaps minimising the need for students to exercise or rely upon their own GenAI literacy skills explicitly. This proactive scaffolding likely enabled both high and low GenAI literacy students to allocate their cognitive resources more effectively during task completion, thus levelling performance differences related to GenAI literacy. Students in the proactive condition benefitted from structured support, reducing the cognitive demands associated with engaging directly with GenAI tools, and enabling those with lower initial GenAI literacy to perform comparably to students with higher sophistication. Previous studies suggest scaffolding can enhance learner performance by decreasing unnecessary cognitive load, especially for those less adept with novel technologies 45, 50. Additionally, by providing structured guidance to learners with varying technological familiarity, this scaffolding approach may help mitigate the risk of a digital divide 8. However, it is also possible that the proactive condition limited opportunities for learners to actively practice and further develop their GenAI literacy skills, a consideration that warrants future investigation. Consequently, educators should balance structured support from GenAI chatbots with opportunities for independent inquiry, encouraging active and strategic engagement that leverages learners’ pre-existing technological literacy.
5.3 Predictive Role of GenAI Literacy in Multimodal Writing (RQ3)
Ordinal logistic regression (RQ3) analyses further substantiated that GenAI literacy significantly predicted students’ independent writing outcomes. Even after controlling for domain-specific content knowledge, initial AI-assisted writing scores, and chatbot interaction conditions, our analyses indicated GenAI literacy uniquely predicted stronger independent writing outcomes in visual data integration, critical thinking, and overall composite scores post-scaffolding removal. By contrast, domain-specific knowledge and chatbot conditions did not significantly predict independent performance outcomes when considered alongside GenAI literacy. These findings reinforce recent arguments positioning digital tool proficiency (especially with GenAI) as a distinct skill set separate from traditional academic knowledge bases 28, 62, 58. Our findings suggest that training and instruction related explicitly to GenAI literacy may effectively support students in independently performing multimodal, complex writing tasks. This unique predictive influence prompts reconsideration of GenAI literacy as a key pedagogical target, warranting purposeful curriculum inclusion rather than simple assumptions of passive skill acquisition.
5.4 Implications for Research and Practice
Our findings have several implications for educational research related to the integration of GenAI in academic writing contexts. First, the identified relationships between GenAI literacy and independent multimodal writing performance highlight the importance of incorporating systematic instruction and explicit assessment of students’ technology-related literacy skills within contemporary curricula. Researchers emphasise that students do not naturally or uniformly acquire sophisticated literacy skills required to critically engage with or use advanced AI technologies in academic contexts 28, 45. Therefore, educators and curriculum designers should proactively embed training on critical aspects of GenAI literacy, such as effectively formulating prompts, interpreting and evaluating AI-generated outputs, and iteratively refining generated content in complex multimodal academic tasks. Future researchers might also explore longitudinal trajectories of students’ GenAI literacy development and potential transfer effects of such skills across various academic domains and types of academic tasks 58, 30. Additionally, careful studies into the cognitive and metacognitive strategies students employ in interaction with GenAI could yield valuable pedagogical insights for developing effective human-AI collaboration models 67, 10.
At a practical level, our findings underscore the importance of balancing external GenAI scaffolding support with opportunities for independent exploration and inquiry. Specifically, passive or reactive chatbot interventions appear more directly beneficial to learners with high GenAI literacy, as such interaction strategies require greater student initiative, metacognitive monitoring, and strategic engagement with the technology 16. In instructional practice, educators might strategically incorporate AI-assisted tools with fading levels of scaffolding to progressively allow students to take greater control over their learning processes, aligning with established theoretical foundations (e.g., Zone of Proximal Development) and scaffolding frameworks 56, 57. Practitioners should also deliberately consider design characteristics and interaction patterns of instructional chatbots, ensuring that chatbot designs offer sufficient flexibility while fostering productive human-AI interactions and learner autonomy 45, 63, 60. Taken together, these implications highlight the need for intentional, critical, and well-informed use of GenAI tools in educational practice, enabling learners to utilise these powerful tools effectively while continually strengthening their independent scholarly skills and cognitive competencies.
5.5 Limitations and Future Directions
Several limitations warrant attention when interpreting the present findings. First, the current study assessed students’ writing performance solely at two temporally adjacent points (with and immediately after the removal of AI scaffolding) within the same experimental session, limiting our ability to make broader claims about delayed or long-term improvements. Future longitudinal studies involving follow-up measures conducted at various time intervals can more definitively establish delayed effects or longer-term skill consolidation. Second, the task’s particular context, healthcare simulations involving multimodal academic writing and visual data interpretation, limits direct generalisability across disciplines or other academic writing types. Replicating or expanding this approach to diverse fields (such as STEM, social sciences, and the humanities) would clarify the robustness and generalisability of current findings. Third, although we distinguished passive from proactive chatbot conditions, additional chatbot features (e.g., specific scaffolding strategies, transparency, adaptivity, fairness) likely shape interactions in complex ways, warranting follow-up studies that evaluate nuanced chatbot design variables. Fourth, qualitative evidence clarifying students’ cognitive processes, metacognitive reflections, and explicit interactions with GenAI tools was beyond our current analysis scope. Future research integrating qualitative approaches (think-aloud protocols, interviews, reflection statements) or detailed log-file analyses of user-AI interactions could richly explain underlying adjustments or strategic variations driven by differing GenAI literacy levels.
6 Conclusion
As GenAI technologies like advanced chatbots become prevalent in education, questions arise regarding their impact on independent thought and student performance. Our study provides empirical evidence that GenAI literacy positively predicts students’ multimodal academic writing performance even after the removal of AI support. Significantly, these GenAI literacy skills influenced outcomes beyond domain-specific knowledge and prior academic performance, highlighting their distinct importance. Additionally, interaction style mattered; passive chatbot support effectively promoted autonomous skill application among AI-literate learners. These results underscore the importance of explicitly nurturing GenAI literacy within curricula and thoughtfully balancing AI assistance with learner autonomy. Future longitudinal studies should explore lasting impacts, informing the responsible integration of GenAI into educational settings.
*Acknowledgments This research was in part supported by the Australian Research Council (DP220101209, DP240100069). L.Y.’s work is fully funded by the Digital Health CRC (Cooperative Research Centre). D.G.’s work was, in part, supported by the DHCRC and Defense Advanced Research Projects Agency (DARPA) through the Knowledge Management at Speed and Scale (KMASS) program (HR0011-22-2-0047). The DHCRC is established and supported under the Australian Government’s Cooperative Research Centres Program. The U.S. Government is authorised to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S. Government.
*Conflict of interest
The authors declare no potential conflict of interest.
*Ethics Statement Ethics approval was obtained from Monash University (Project ID: 37307). Informed written consent was also obtained from students before the data collection.
*Data Availability
For any queries regarding the data and analysis codes, please contact the corresponding author.