### v2

# The Agency Gap: How Generative AI Literacy Shapes Independent Writing after AI Support

Abstract
Generative AI (GenAI) tools are rapidly transforming higher education, yet little is known about how students’ GenAI literacy shapes their ability to perform independently once such support is removed. This study investigates what we term the agency gap, introduced as the extent to which GenAI literacy predicts student writing performance in contexts that require self-initiation and regulation. Seventy-nine medical and nursing students completed multimodal academic writing tasks based on visual data, supported either by a reactive or proactive GenAI chatbot, followed by a parallel task without AI support. Writing was evaluated across insightfulness, visual data integration, organisation, linguistic quality, and critical thinking. Results showed that GenAI literacy predicted independent writing performance only in the reactive condition, where students had to actively mobilise their own strategies. Mediation analyses revealed no indirect effect via in-task performance, indicating that GenAI literacy acts as a direct, task-general competence rather than a proxy for domain knowledge or other literacies. By contrast, proactive scaffolding equalised outcomes across literacy levels, reducing reliance on learners’ GenAI literacy. The agency gap highlights when GenAI literacy matters most, with implications for designing equitable AI-supported learning environments that both leverage and mitigate differences in students’ GenAI literacy.
keywords:
[…]
Academic writing is an essential practice in higher education. It helps students develop critical thinking, effective communication, and the ability to deal with complex ideas and information (Lea and Street, 1998; Oshima and Hogue, 2007). Students need to interpret and synthesise data, create logical arguments, and assess different perspectives (Durst, 1987; Paltridge, 2004). Increasingly, academic writing also involves multimodal literacy, requiring students to accurately interpret, synthesise, and incorporate varied visual data sources, such as graphs, charts, or images, into their arguments. This skill reflects the increasingly multimodal nature of contemporary academic discourse and demands additional capabilities concerning visual-data integration (Lea and Street, 2006; Oshima and Hogue, 2007; Jewitt, 2006). Academic writing is crucial for both academic success and professional competency (Swales and Feak, 1994; Wingate, 2012). Employers value strong academic writing skills as they help individuals write persuasive reports, create coherent presentations, and solve problems effectively (Council, 2012). Yet, academic writing is linked to cognitive and metacognitive processes. It encourages students to monitor and control their thought processes, facilitating metacognitive awareness and self-regulation (Flavell, 1979). This engagement is essential for students’ ability to plan, evaluate, and reflect on their learning strategies, leading to a deeper understanding and better retention of information (Hacker et al., 2009).
As technology advances, academic writing can evolve from being solely an individual process to one supported by various tools (Nguyen et al., 2024). With the increasing availability of generative AI (GenAI) tools (e.g., ChatGPT and Gemini), students have more opportunities to enhance their academic writing performance (Yan, 2023; Bašić et al., 2023). Using GenAI tools in writing is becoming more common (Kim et al., 2025), making it urgent to understand how these external supports impact students’ skill development (Liu et al., 2024). GenAI tools can provide feedback, generate content ideas, or assist in structuring arguments, offering new ways for students to engage with writing tasks (Cardon et al., 2023). Yet their growing use raises a critical question: do GenAI tools genuinely build students’ independent writing skills, or do they merely provide temporary performance boosts (Fan et al., 2025; Yan et al., 2024a)? This issue is especially salient in multimodal contexts, where students are required to write while interpret and synthesise visual information (Liu et al., 2024). While research has begun to explore how GenAI assists writing (Nguyen et al., 2024), little is known about whether its benefits transfer once the technology is removed, or how different interaction strategies, such as reactive replying versus proactive scaffolding, shape that transfer (Jin et al., 2025).
A further underexplored dimension is the role of students’ GenAI literacy. Beyond general academic or domain-specific literacies, GenAI literacy encompasses knowledge of how AI systems operate, the ability to formulate effective prompts, the capacity to critically evaluate outputs, and an awareness of associated risks and ethical considerations (Jin et al., 2024; Bozkurt, 2024; Ng et al., 2021). Students with higher levels of GenAI literacy may use GenAI tools more effectively while also sustain their performance once such support is removed. Yet it remains unclear whether GenAI literacy acts indirectly, by improving AI-assisted work that then transfers, or directly, as a transferable competence that enables independent performance.
In this study, we investigate what we term the agency gap—the extent to which GenAI literacy predicts independent academic writing when learners must self-initiate and regulate their engagement with AI. Seventy-nine higher education students completed multimodal academic writing tasks with either reactive (responding only to queries) or proactive (actively scaffolding) GenAI support, followed by a parallel task without AI. Results showed that GenAI literacy predicted independent writing performance only in the reactive condition, where learner agency was essential, and not in the proactive condition, where scaffolding equalised outcomes across literacy levels. Mediation analyses further revealed no indirect effect via AI-supported performance, indicating that GenAI literacy acts as a direct, task-general competence rather than a proxy for domain knowledge or writing proficiency. By conceptualising and evidencing the agency gap, this study advances understanding of when and how GenAI literacy matters and highlights two complementary levers for practice: cultivating GenAI literacy to support independent performance, and designing proactive scaffolds to mitigate disparities when learner agency cannot be assumed.
2 Background
2.1 GenAI and Academic Writing
Academic writing is a core skill in higher education, requiring students to synthesise ideas effectively, engage critically with evidence, and communicate coherently. At its heart, strong writing demonstrates insightfulness, where students provide original and nuanced interpretations by deeply analysing subject matter and connecting ideas across sources (Kuhn, 2005; Wingate, 2012). Contemporary academic discourse also increasingly requires visual data integration, as students are increasingly expected to interpret and incorporate charts, diagrams, and other graphical information into their arguments (Lea and Street, 2006; Oshima and Hogue, 2007; Jewitt, 2006). Equally essential is organisation and coherence, ensuring that arguments flow logically and ideas are structured in ways that support reader comprehension (Hyland, 2021; Swales and Feak, 1994). The quality of academic writing further relies on linguistic quality, reflected in precise grammar, appropriate vocabulary, and adherence to academic style conventions (Biber and Gray, 2010; Hyland and Milton, 1997). Finally, rigorous critical thinking underpins effective writing by enabling students to evaluate evidence, consider alternative perspectives, and develop reasoned arguments (Moon, 2007; Facione, 2000). A holistic assessment of academic writing, therefore, demands comprehensive evaluation across all of the above dimensions, to capture the multifaceted nature and reflect the complexity student face in authentic academic writing.
More recently, the advancements in GenAI have transformed writing tools into sophisticated assistants for diverse writing tasks (Lee et al., 2024). Research has produced systems for AI-powered writing support, such as CoAuthor, which provides sentence-level suggestions integrated into the writing process (Lee et al., 2022), and tools that generate text continuations based on user input (Jakesch et al., 2023). Studies show that students expect GenAI to serve multiple roles, including as an assistant, tutor, and collaborator (Kim et al., 2025), though they also have concerns about its ethical and practical implications (Wang, 2025). Further work has explored the behavioural patterns that emerge during GenAI-assisted writing (Shibani et al., 2023; Nguyen et al., 2024), finding that iterative and interactive engagement correlates with improved outcomes (Nguyen et al., 2024) and that engagement in higher-order cognitive processes is linked to greater textual sophistication (Yang et al., 2025). Despite these advances, current research primarily addresses text-based writing, leaving GenAI’s role in multimodal writing that involves visual data integration largely unexplored. Furthermore, the relationships between students’ GenAI literacy and their subsequent writing performance remain under-investigated, limiting the ability to offer tailored, literacy-specific support. Finally, little research has examined how different chatbot interaction strategies, such as reactively responding versus proactively scaffolding, impact academic writing. This is a critical gap, as interaction strategies significantly shape student engagement and learning outcomes during human-AI collaboration (Nguyen et al., 2024; Jin et al., 2025).
2.2 GenAI Chatbot Interaction Strategies
The integration of GenAI chatbots, particularly conversational chatbot, into educational technologies has shown a growing trend as a solution to offer scalable and dynamic support to empower existing educational technologies (Akpan et al., 2024; Bozkurt, 2023). Conversational chatbots are designed to infuse interactivity into learning environments, providing explanations, clarifications and insights that enhance students’ cognitive and emotional engagement (Wu et al., 2025). The most widely adopted form of GenAI chatbot in education is the conventional reactive chatbot. These chatbots behave in a reactive manner, responding to student-initiated queries by delivering contextually relevant information. Their value lies in providing clarity and explanation when prompted, which can effectively support student understanding in knowledge-intensive tasks. Empirical studies have demonstrated the use of reactive chatbots can enhance learning across varied contexts and populations. They have been found to support knowledge-based dialogue and boost creative problem-solving performance among postgraduate students (Song et al., 2025), increase motivation and reduce stress while enabling self-paced learning in secondary programming education (Ait Baha et al., 2023), and improve science knowledge, engagement, and motivation when compared with rule-based systems (Ng et al., 2024). However, despite these benefits, reactive chatbots require students equipped with ability to initiate effective queries, meaning that learners with limited GenAI literacy may struggle to maximise their value.
Recent work has moved beyond the limitations of reactive chatbots toward the design of proactive scaffolding chatbots informed by educational theory. Scaffolding, rooted in Vygotsky’s zone of proximal development, supports learners by breaking complex tasks into smaller steps, asking guiding questions, and providing tailored feedback that promotes self-regulation (Gibbons, 2002). By operationalising these principles, proactive chatbots engage learners through structured narratives, dynamically adapt to student input, and reduce reliance on prompting skills, thereby providing more equitable support. Empirical studies have demonstrated their effectiveness across diverse contexts. For instance, they have improved comprehension of complex visualisations by guiding step-by-step interpretation (Yan et al., 2025), enhanced instructional design performance through structured exploration and feedback (Bai et al., 2024), facilitate personalised learning through structured tutor–student dialogues that enhance knowledge acquisition and critical analysis (Looi and Jia, 2025), and support self-regulated strategy development in reading through interactive guidance that enhanced metacognitive and motivational engagement (Pan et al., 2025). However, there is still limited research on how different chatbot interaction strategies shape students’ writing performance, particularly in relation to differences in their levels of GenAI literacy.
2.3 Generative AI Literacy
Given the increasing use of GenAI in educational contexts, particularly in academic writing, it has become essential to investigate and consider students’ GenAI literacy when exploring AI-assisted writing practices. While general AI literacy has been defined and explored extensively in previous frameworks (Long and Magerko, 2020; Ng et al., 2021), the unique characteristics of generative technologies necessitate a more specialised literacy framework. Researchers have highlighted that GenAI literacy involves not only understanding how generative tools function, but also mastering practical skills to effectively prompt these tools, critically evaluate AI-generated content, and navigate ethical and contextual complexities specific to GenAI-generated outputs (Zhao et al., 2024; Bozkurt, 2024). Despite recognition of the importance of GenAI literacy for meaningful interaction with tools like ChatGPT and Gemini, there remains limited empirical research that explicitly examines how students’ varying levels of GenAI literacy impact their academic writing performances when GenAI assistance is provided or withdrawn (Annapureddy et al., 2024; Lyu et al., 2024). Addressing this gap is imperative, as understanding the relationship between students’ GenAI literacy and their academic writing outcomes can inform practical strategies for educators seeking to effectively integrate these advanced technologies into multimodal academic writing instruction (Chiu, 2024; Annapureddy et al., 2024).
2.4 Domain Knowledge
Domain knowledge is an essential consideration when investigating AI-assisted multimodal academic writing, particularly when visual information and interpretation are integral to the task. Students’ capacity to successfully integrate visual data into meaningful written discussions depends heavily on their familiarity with relevant subject areas (Alexander, 1992; Tricot and Sweller, 2014). This underscores the necessity of evaluating and accounting for the domain knowledge possessed by individual learners, especially when examining the relationship between their GenAI literacy and writing performance (Alexander and Judy, 1988). Without considering domain proficiency, researchers may inadvertently attribute students’ challenges in producing coherent multimodal arguments or effectively utilising GenAI supports to deficiencies in GenAI literacy alone, rather than identifying underlying gaps in subject-specific knowledge (Alexander, 1992). Accordingly, research into GenAI literacy for multimodal academic writing contexts should adopt a nuanced approach, explicitly acknowledging the role of domain-specific expertise alongside GenAI literacy.
2.5 Research Gaps
The current study aimed to address the aforementioned gaps by examining how students’ GenAI literacy affects their academic writing performance with and without GenAI support. Specifically, we investigated whether these effects differ based on chatbot interaction strategies (reactive vs. proactive) and account for students’ domain-specific knowledge to ensure an accurate interpretation of the effects of GenAI literacy. The following research questions were investigated:
RQ1: To what extent does students’ GenAI literacy relate to their academic writing performance, after controlling for their visualisation literacy, and how does this relationship differ across chatbot interaction types (reactive vs. proactive) and at the AI-support versus AI-removal stages?
RQ2: Do students with low versus high GenAI literacy differ significantly in their academic writing performance at the AI-support and AI-removal stages, and do these differences differ based on chatbot interaction type (reactive vs. proactive)?
RQ3: To what extent does students’ GenAI literacy predict their academic writing performance at the AI-removal stage after controlling for domain-specific knowledge, chatbot interaction type (reactive vs. proactive), and writing performance at the AI-support stage?
RQ4: To what extent does students’ academic writing performance during the AI-support stage mediate the relationship between their GenAI literacy and their performance at the AI-removal stage?
3 Methods
3.1 Learning Contexts
The learning context involved writing based on visual information, a critical skill in academic writing, closely aligned with tasks such as IELTS Academic Writing Task 1 (IELTS, 2025). Participants received three distinct visual representations (a bar chart, a network diagram, and a ward map; described in detail in Section 3.3.1). They were instructed to interpret these visuals and compose a short written answer (150 words) addressing a question about team performance (e.g., "How the two nurses managed the primary patient (Amy) while attending to other beds, focusing on their task prioritisation, verbal communication, and stress levels?"). The visuals were specifically designed using data from authentic learning contexts, such as healthcare simulations, ensuring participants had no prior exposure to these materials. Thus, any potential effects stemming from participants’ prior familiarity with the presented information were minimal. The study was conducted through an online environment (details in Section 3.3.2).
3.2 Participants
A total of 81 participants completed the study. However, two participants were excluded from the analysis due to providing written responses shorter than 20 words, resulting in a final sample of 79 participants. This exceeded the minimum sample size (72 participants; 36 per condition) required by a priori power analysis (effect size , power = 0.80, ). Participants were recruited via Prolific111https://www.prolific.co/ and compensated £8 for approximately one hour of participation. Participants were randomly assigned to one of two experimental conditions: reactive chatbot (39) or proactive chatbot (40). All participants were current students or graduates from medical or nursing programs. The sample comprised 45 females and 34 males, primarily aged 18-24 (28) and 25-34 (31); fewer participants belonged to age groups 35-44 (11), 45-54 (6), 55-64 (2), and 65 or older (1). Participants were distributed geographically across North/Central America (34), Europe (21), Africa (15), Australia (4), South America (2), and other regions (3). Educational backgrounds included Bachelor’s degrees (36), high school diploma or equivalent (18), Master’s degrees (11), vocational training or diplomas (7), Doctoral degrees (4), and other qualifications (3). Familiarity with GenAI tools was predominantly intermediate (42), beginner (25), advanced (9), expert (2), and none (1).
3.3 Materials and Apparatus
3.3.1 Learning Materials
The three visual information representations used for learning were generated from data collected during a high-fidelity healthcare simulation. In this simulation, students worked in teams of four (comprising two primary and two secondary nurses) to navigate acute clinical scenarios within a realistic, ward-like environment. The simulated setting included hospital beds, medical equipment, and manikins designed to replicate patient vital signs. The objective of the simulation was to enhance students’ prioritisation skills, communication strategies, and collaborative teamwork by replicating realistic clinical emergencies. Throughout the simulation, students’ positional information (x-y coordinates and body orientation) was captured using an indoor positioning system; audio interactions were recorded through wireless headset microphones; and physiological data, specifically heart rates, were collected using Fitbit Sense wristbands.
The multimodal data were used to generate three visuals (Figure 1), each with progressively increasing complexity. First, a bar chart summarises time spent on different behaviours derived from positional data. This visual information illustrates team-level resource allocation and prioritisation strategies [Anonymised]. Second, a communication network diagram combining positional and audio data to reveal the frequency and direction of verbal interactions among nurses, the doctor, the patient, and the patient’s relative, thereby depicting participant roles and their level of involvement [Anonymised]. Third, a ward map integrates positional, audio, and physiological data. Inspired by visual and spatial analytics methods used in professional sports (Goldsberry, 2012), the ward map utilises heatmap visualisation techniques to show communication intensity, trace spatial movements, and highlight locations of peak physiological arousal via heart rate data.
3.3.2 Research Platform
The academic writing tasks were conducted using a custom-built platform comprising three main components. The first component, the display area (Figure 2), presented one visual analytic at a time to minimise visual clutter and prevent information overload (Ellis and Dix, 2007). Participants could freely navigate between the three visual information representations using the navigation arrows positioned on either side of the component. The second component, the writing space (Figure 2), provided participants with an area to articulate their analytical insights by synthesising information presented in the visual information. The third component varied based on the experimental condition. In the AI-supported conditions, participants had access to a chat function (Figure 2), offering real-time support from GenAI agents, which responded either reactively or proactively. Conversely, in conditions without AI support, the chat function was replaced by an instruction component providing standard task instruction without AI assistance. The overall platform design emphasised simplicity (Stone et al., 2005), enabling participants to focus primarily on comprehending the visual information and effectively completing the tasks without being overwhelmed by complex functionalities.
3.3.3 Chatbot Interaction Design
We developed two distinct GenAI chatbots, namely the reactive and the proactive, primarily differing in their interaction types and learner-support strategies (Figure 3). The purpose behind these design choices was to explore how interaction type associates with the writing performance of learners with varying GenAI literacy levels. While both chatbots introduced users with a welcome message explaining the writing task, the reactive chatbot provided a description of its reactive and supportive features and responds only to users’ questions without following up with further interactions (Ma et al., 2023; Yan et al., 2024b). In contrast, the proactive chatbot actively initiated conversations with users and continually engaged them through domain expert-designed questions and structured narrative pathways (see Appendix D) (Park et al., 2023). Rather than simply reacting, the proactive chatbot offered structured questioning sequences and iterative feedback, progressively enhancing learners’ understanding of visual information based on Hattie and Timperley (2007)’s feedback model (e.g., task and process-level). For example, when guiding learners through the ward map analyses, the proactive chatbot first prompted them towards interpreting the distribution of hexagons in the ward (positional data). Once it confirmed that the user had clearly understood this aspect, it then moved onto the next layer of interpretation, specifically the intensity of the colour (verbal communication). This scaffolded questioning approach aligned closely with established conceptual guidelines for interpreting visual information (Xun and Land, 2004), systematically aiming to enhance learners’ critical interpretation skills.
Building upon the VizChat prototype (Yan et al., 2024b), both chatbots adopted an advanced agentic architecture composed of four interconnected modules. Prompts were categorised into reactive and proactive types corresponding to the aforementioned interaction strategies. At its core, the GenAI agent managed and synthesised inputs from visual information and user prompts, utilising Retrieval-Augmented Generation (RAG) to efficiently retrieve data-informed content from its knowledge database (Gao et al., 2023). This comprehensive knowledge repository contained contextual information, visualisation details, and interaction logs that enabled personalised and relevant interactions. Finally, synthesised prompts were processed by the multimodal GPT-4o AI model, generating coherent and context-aware responses that enriched the database and progressively improved the quality of subsequent learner interactions.
3.4 Measures and Instruments
3.4.1 Generative AI Measurement (GLAT)
We assessed learners’ generative AI literacy using the Generative AI Literacy Assessment Test (GLAT) (Jin et al., 2024). This 20-item instrument was selected for its established validity and reliability. Face validity was ensured through expert review and revision, while content validity was evaluated with 204 participants to confirm item relevance, comprehensiveness and comprehensibility. Structural validity and internal consistency were confirmed using classical test theory and item response theory, supporting a reliable two-parameter logistic (2PL) model (Cronbach’s ) and robust factor structure (, , , 90% CI ). As the 2PL model accounts only for item difficulty and discrimination, we applied a correction-for-guessing formula (Thorndike et al., 1991) to control for potential guessing effects (Eq. 1).
In this formula, represents the final score corrected for guessing behaviours, denotes the number of correctly answered items by a given participant, denotes the number of wrongly answered items, and denotes the number of answer choices per item, which is four for the GLAT. The full test is available in Appendix A.
3.4.2 Domain Knowledge Measurement (VLAT)
Visualisation literacy constitutes essential domain knowledge for effective multimodal writing (Felten, 2008; Lee et al., 2016). We assessed participants’ domain knowledge in visualisation literacy using the Visualisation Literacy Assessment Test (mini-VLAT) (Pandey and Ottley, 2023). This 12-item instrument was selected for its established validity and reliability, which are comparable to those of the original 53-item VLAT (Lee et al., 2016), while providing a briefer assessment suitable for our extended study. Its concise structure helped minimise participants’ cognitive load. Structural validity and internal consistency for this instrument have previously been confirmed (Pandey and Ottley, 2023). To account for potential guessing effects, we also applied the correction-for-guessing formula (Thorndike et al., 1991) to participants’ scores. The types of visualisations and their corresponding questions are provided in Appendix B.
3.4.3 Academic Writing Assessment Rubric
A comprehensive assessment rubric was developed to evaluate the participants’ academic writing performances in tasks involving visual data analyses. Based on prior literature and established theoretical grounding within academic writing research (as elaborated in Section 2.1), we identified five distinct evaluation metrics (Table 1): Insightfulness,Visual Data Integration, Organisation and Coherence, Linguistic Quality, and Critical Thinking. Each criterion was evaluated using a 5-point Likert scale (Level 1–Poor to Level 5–Excellent), with comprehensive descriptors and examples available in the full rubric (see Appendix C).
| Metric | Description | Reliability () |
| Insightfulness | Assesses the originality, depth, and relevance of insights. High-quality responses identify subtle patterns, offer holistic interpretations, and provide conclusions that extend beyond obvious points. Low-quality responses are surface-level observations or simplistic interpretations (Kuhn, 2005; Chanock, 2000). | 0.82 |
| Visual Data Integration | Measures the effectiveness and clarity with which visual data (charts, graphs, communication networks, heart rate information) is explicitly interpreted, combined, and connected into the written analysis. High-quality responses synthesize multiple visual data sources cohesively, while low-quality responses merely mention data without synthesizing or miss integrating data entirely (Jewitt, 2006). | 0.85 |
| Organisation and Coherence | Evaluates the clarity, logical structure, and connectedness of ideas presented in the writing. High-quality responses have logical structure, smooth transitions between points, and clearly organised arguments. Low-quality responses lack clarity, exhibit abrupt transitions, or are disorganised (Paltridge, 2004). | 0.75 |
| Linguistic Quality | Reflects grammar accuracy, vocabulary appropriateness, sentence fluency, and overall clarity of written expression. High-quality responses utilize clear, precise, and varied language with minimal errors, while low-quality responses contain grammatical errors, limited vocabulary, or unclear expression (Biber and Gray, 2010). | 0.89 |
| Critical Thinking | Evaluates the level of careful thought, complexity, consideration of alternative viewpoints, and strength of conclusions and recommendations. High-quality responses deeply analyze complexities and implications, considering broader contexts, contingencies, or solutions. Low-quality responses show limited or superficial analytical engagement and fail to address alternative perspectives or implications (Facione, 2000; Moon, 2007). | 0.80 |
3.4.4 Scoring Procedures and Inter-Rater Reliability
The scoring process followed detailed guidelines based on the rubric provided in Appendix C. Two independent, trained coders rated each participant’s writing samples at each study phase (AI-support and AI-removal). Coders were trained extensively on the rubric definitions through practice scoring sessions with sample responses (n=10), followed by discussion and calibration of understanding to maximise rating consistency. All essays were then independently blind-rated by both coders. After ratings were fully completed, reliability analyses were performed (Spearman’s correlations), demonstrating high inter-rater reliability ( in all cases; Table 1). Final participant scores for each dimension, at each stage, were averaged across the two coders, producing a single composite rating for each dimension and subsequently an overall Composite Score. To ensure transparency and replicability, example excerpts representing each scoring level were provided in the detailed rubric (see Appendix C).
3.5 Study Procedure
We conducted a two-phase, mixed-method experimental study to examine how GenAI literacy and AI interaction type (reactive vs. proactive chatbot) influenced participants’ academic writing performance with visual information. Participants were recruited via Prolific and directed to a Qualtrics222https://www.qualtrics.com/ survey, where they provided informed consent, demographics, and completed literacy assessments (GLAT and mini-VLAT). After reviewing contextual information on the healthcare simulation (Section 3.3.1), they were randomly assigned to either the reactive or proactive chatbot assistance condition. In the AI-supported phase, participants composed a 150-word written response interpreting the provided multimodal visuals with real-time chatbot support. Subsequently, in the AI-removal phase, participants completed a similar task with new visual data from a different simulation scenario without chatbot access (Figure 4). The study lasted approximately one hour. Ethical approval was obtained from [Anonymised] University (Project Number: Anonymised).
3.6 Analysis
3.6.1 Correlation Analysis (RQ1)
To address RQ1, we conducted a series of correlation analyses to investigate the relationships between students’ GenAI literacy and their multimodal academic writing performance. The analysis proceeded in two stages.
First, we examined the direct associations between literacy scores and writing outcomes. We explored the relationships of both GenAI literacy (as measured by the GLAT) and visualisation literacy (as measured by the mini-VLAT), which served as a proxy for the specific domain knowledge required for the task, with writing performance scores. Prior to analysis, Shapiro-Wilk tests were conducted on all variables, which indicated that the literacy and writing scores were not normally distributed (). Consequently, we employed Spearman’s rank-order correlation () for all subsequent analyses. These correlations were computed separately for the reactive and proactive chatbot conditions at both the AI-support and AI-removal time points. The analysis was applied to the composite writing score as well as to each individual rubric dimension, including insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking.
To provide a more rigorous test of our research question, the second stage of our analysis aimed to isolate the unique contribution of GenAI literacy. We performed a partial correlation analysis to isolate the influence of GenAI literacy from that of task-specific domain knowledge (i.e., visualisation literacy), thereby controlling for potential confounding effects. This involved first computing the zero-order Spearman’s correlation between GLAT and VLAT scores to assess their degree of overlap within each condition. Subsequently, we conducted a semi-partial correlation to determine the association between GenAI literacy and writing performance after statistically removing the variance attributable to visualisation literacy. This conservative approach allowed for a more precise estimation of the unique relationship between GenAI literacy and writing performance. This analysis was conducted consistently with the primary correlations, examining each chatbot condition and time point separately. For all analyses, statistical significance was set at .
3.6.2 Comparative Analysis (RQ2)
To investigate whether students with varying levels of GenAI literacy performed differently in academic writing (RQ2), we conducted a series of comparative analyses. As a preliminary step, we first examined whether domain knowledge was a significant confounding variable in relation to writing performance. To do this, we examined if differences in visualisation literacy were associated with writing performance. Participants in each condition were divided into low and high visualisation literacy groups based on a median split of their VLAT scores. We chose the median split procedure to ensure balanced group sizes and improve interpretability. Mann-Whitney tests were then performed to compare the writing scores between these groups, which are appropriate for the ordinal nature of the rubric scales and the non-normal distribution of the data (verified through Shapiro-Wilk tests). This initial analysis allowed us to determine whether domain knowledge was a significant factor in its own right before proceeding to our primary question.
Having established the role of domain knowledge, we conducted the main analysis for RQ2, to investigate if students with varying level of GenAI literacy performed differently in academic writing. Participants were again categorised into low and high literacy groups, this time based on a median split of their GLAT scores within each condition. Mann-Whitney tests were also selected for this analysis.
For both analyses of the VLAT and GLAT scores, tests were conducted separately for each rubric dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking) as well as for the composite score, across both the AI-support and AI-removal stages. Statistical significance was set at and adjusted for multiple comparisons using the Bonferroni method. Significant differences were further interpreted using effect sizes calculated as the Rank-Biserial correlation.
3.6.3 Regression Analysis (RQ3)
To investigate the extent GenAI literacy predicted writing performance after controlling for domain-specific knowledge (RQ3), ordinal logistic regression analyses were conducted. As shown in the following formula (2), the dependent variables for regression analyses were participants’ academic writing performance scores during the AI-removal period for each dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) separately, as well as the composite writing score. Each outcome variable, measured using ordinal scales (; scores from 1 to 5; refers to each ordinal threshold, ), justified the use of ordinal logistic regression rather than linear regression. Primary predictor variables included standardised GenAI literacy scores (), domain knowledge (; VLAT score), chatbot intervention type (; binary-coded: 0 = reactive chatbot, 1 = proactive chatbot), and participants’ AI-assisted scores (; to control writing status during chatbot support availability). Additionally, denotes the threshold/intercept parameter for each category boundary, and , and represent regression coefficients for the predictors.
Prior to analysis, GenAI literacy scores were standardised (M = 0, SD = 1) to facilitate regression coefficient interpretation. Ordinal logistic models were fitted using a logit link function and estimated with Maximum Likelihood Estimation via the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm in the Python Statsmodels library. Before analyses, multicollinearity was examined through variance inflation factor (VIF) values, and proportional odds assumptions were evaluated using the Brant test. Statistical significance of model predictors was evaluated at the level. Regression results were interpreted by examining regression coefficients (), standard errors (SE), test statistics (z-values), significance levels (-values), and odds ratios (ORs). Odds ratios greater than one indicated an increased likelihood of scoring higher on the writing assessment due to higher levels of GenAI literacy, whereas odds ratios less than one indicated a decreased likelihood.
3.6.4 Mediation Analysis (RQ4)
To investigate the mechanism through which GenAI literacy influences independent writing performance (RQ4), we conducted a mediation analysis. The primary hypothesis was that the relationship between a student’s GenAI literacy (the independent variable, X) and their subsequent independent writing performance at the AI-removal stage (the outcome variable, Y) would be mediated by their writing performance during the AI-support stage (the mediator, M). This analysis was designed to test whether GenAI literacy enhances in-task performance, which in turn transfers to improved independent writing capabilities. The mediation analysis was specified using two linear models. The mediation model (Eq. 3) estimated the effect of GenAI literacy on AI-support performance, while the outcome model (Eq. 4) estimated the effect of both GenAI literacy and AI-support performance on independent writing performance. Here, captures the effect of GenAI literacy on AI-support performance, captures the effect of AI-support performance on independent writing performance controlling for GenAI literacy, and represents the direct effect of GenAI literacy on independent writing performance. The indirect effect is given by .
We employed a modern bootstrapping procedure with 5,000 resamples to estimate the indirect effect () and its 95% confidence interval. This method was selected for its statistical power and because it does not assume a normal sampling distribution for the indirect effect, making it more robust than traditional tests. The analysis was performed separately for the reactive and proactive chatbot conditions to examine whether the mediational pathway differed based on the interaction strategy. A significant indirect effect was determined if the 95% confidence interval did not contain zero.
4 Results
4.1 Correlation Between GenAI Literacy, Domain Knowledge, and Writing Performance (RQ1)
Visualisation Literacy and Writing Performance
Spearman’s rank-order correlations between participants’ VLAT scores and writing performance across each rubric and composite score, separated by chatbot condition and time point, indicated no statistically significant associations (Figure 5). In the reactive condition (), correlations ranged from to , all , with the largest being Organisation and Coherence at AI-removal stage (). In the proactive condition (), coefficients ranged from to , again none of the coefficients was significant. These results suggest that domain-specific visualisation literacy did not significantly associate multimodal academic writing performance.
GenAI Literacy and Writing Performance
Spearman’s rank-order correlations revealed that GenAI literacy was positively associated with writing performance, particularly in the reactive chatbot condition during the AI-removal stage (Figure 6). Significant correlations were observed with Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and the Composite Score (). A smaller but significant effect was also found for Linguistic Quality at the AI-support stage (). In contrast, no significant correlations were observed in the proactive condition.
Partial Correlation Analysis Given the moderate overlap between visualisation literacy and GenAI literacy (reactive: ; Proactive: ), we conducted partial correlations controlling for VLAT scores to rule out confounding effects. The results confirmed that the relationships between GenAI literacy and writing performance remained robust even after controlling for visualisation literacy, particularly in the reactive condition during the AI-removal stage. All significant correlations observed in the zero-order analysis maintained statistical significance in the partial correlation analysis. Specifically, in the reactive condition during AI-removal, partial correlations (controlling for VLAT) were: Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and Composite Score (). Notably, the partial correlation coefficients were similar to the zero-order correlations, with differences ranging from -0.04 to +0.05, suggesting that the observed relationships between GenAI literacy and writing performance are largely independent of visualisation literacy. In contrast, the proactive condition showed no significant partial correlations at either time point, consistent with the zero-order correlation findings (Figure 7).
4.2 Comparisons Between Low and High GenAI Literacy Groups (RQ2)
The Mann-Whitney U tests examining visualisation literacy group differences revealed no statistically significant differences between high and low visualisation literacy groups in either condition across any of the measured writing performance metrics or time points (all ). In the reactive condition (), participants were divided based on the median VLAT score into low () and high () groups. Mann-Whitney U tests were performed separately for each rubric dimension at both the AI-support and AI-removal stages. While descriptive patterns suggested that high VLAT participants tended to achieve slightly higher median scores on some metrics during the intervention phase (e.g., Insightfulness: Low VLAT Mdn = 2.00, High VLAT Mdn = 2.50; Organization and Coherence: Low VLAT Mdn = 2.00, High VLAT Mdn = 3.00), these differences were not statistically significant. The largest effect size observed was for Organization and Coherence during the post phase (; small effect), but this did not reach conventional significance levels. Similarly, in the proactive condition (low VLAT (), high VLAT (), no significant differences were observed between VLAT groups, with generally small effect sizes and minimal differences in median scores, particularly during the post phase. These findings suggest that within the context of AI-supported multimodal academic writing tasks, visualisation literacy levels did not substantially differentiate writing performance outcomes, regardless of whether participants received reactive or proactive AI support.
Turning to GenAI literacy, participants in the reactive chatbot condition () were grouped via a within-condition median split into low () and high () literacy groups. Medium to large differences emerged primarily at the AI-removal stage (without AI support), with high GenAI literacy participants consistently scoring higher across several writing dimensions. Specifically, at the AI-removal stage, significant differences were observed for Insightfulness (), with the high literacy group achieving higher scores (Mdn = 2.75, IQR = 1.00) compared to the low literacy group (Mdn = 2.00, IQR = 0.00); Visual Data Integration (), with higher scores for the high literacy group (Mdn = 2.50, IQR = 0.50) versus low literacy group (Mdn = 2.00, IQR = 0.00); and Organisation and Coherence (), again favouring high literacy (Mdn = 2.75, IQR = 0.50) over low literacy participants (Mdn = 2.00, IQR = 0.50). Additionally, a significant difference was found for the overall Composite Score (), with higher scores among high literacy (Mdn = 13.25, IQR = 3.38) than low literacy participants (Mdn = 11.00, IQR = 1.50). At the AI-support stage (with GenAI assistance), only Linguistic Quality revealed a significant difference (), with the high literacy group scoring higher (Mdn = 3.00, IQR = 0.50) compared to the low literacy group (Mdn = 3.00, IQR = 0.00). On the other hand, participants in the proactive chatbot condition () were similarly grouped into low () and high () GenAI literacy categories based on median splits of GenAI literacy scores. However, Mann-Whitney U tests revealed no significant differences between the low and high literacy groups for any of the rubric metrics at either the AI-support or AI-removal stages.
4.3 Ordinal Logistic Regression Predicting Writing Performance (RQ3)
For each ordinal logistic regression model, proportional odds assumptions were met (Brant test, all ), and multicollinearity diagnostics indicated no problematic multicollinearity (variance inflation factor for all predictors). Results demonstrated that higher GenAI literacy significantly predicted superior writing performance after AI assistance was withdrawn for three rubric criteria: Visual Data Integration, , , , , ; Critical Thinking, , , , , ; and overall writing quality measured through the Composite Score, , , , , . Specifically, each one-unit increase in standardised GenAI literacy was associated with a 14% greater likelihood of achieving higher ratings in Visual Data Integration, a 15% greater likelihood for Critical Thinking, and an 11% greater likelihood to receive higher overall Composite Score. However, GenAI literacy did not significantly predict scores on the Insightfulness, Organisation and Coherence, and Linguistic Quality dimensions. Additionally, participants’ AI-assisted writing scores significantly predicted their subsequent writing performance across all rubric dimensions: Insightfulness, , , , ; Visual Data Integration, , , , ; Organisation and Coherence, , , , ; Linguistic Quality, , , , ; Critical Thinking, , , , ; and overall Composite Score, , , , . Neither domain-specific knowledge (VLAT scores) nor chatbot interaction type (proactive vs. reactive) significantly predicted writing performance in any regression models. Thus, the predictive relationship of GenAI literacy with writing performance was robust regardless of participants’ domain-specific knowledge or chatbot intervention type.
4.4 Mediation Analysis of GenAI Literacy and Writing Performance (RQ4)
The bootstrap mediation analysis was conducted to determine if the relationship between GenAI literacy and independent writing performance was explained by students’ performance on the AI-supported task. The results, summarised in Table 2, did not support the hypothesised mediational pathway (GenAI Literacy AI-Support Performance Independent Performance) in either the reactive or proactive condition.
In the reactive condition, the analysis first confirmed a significant total effect, establishing an overall positive relationship between higher GenAI literacy and stronger independent writing performance (Path c: , ). However, the results did not support the proposed mediational pathway, as the indirect effect was negligible and not statistically significant (Indirect Effect = 0.032, 95% CI [-0.022, 0.108]). Crucially, after accounting for this non-significant indirect path, a notable direct effect between GenAI literacy and independent performance persisted (Path c’: ). This finding demonstrates that the overall benefit of GenAI literacy in the reactive condition was largely independent of the proposed mechanism.
Similarly, in the proactive condition, the indirect effect was not statistically significant (Indirect Effect = 0.022, 95% CI [-0.055, 0.102]). Consistent with prior analyses, GenAI literacy demonstrated no significant total or direct effects on writing performance within this condition, emphasising that it was not a determining factor for student outcomes when proactive scaffolding was provided.
| Parameter | Reactive | Proactive |
| Paths | ||
| Path a (X M) | 0.069 | 0.051 |
| Path b (M Y | X) | 0.468 | 0.438 |
| Path c’ (Direct Effect) | 0.140 | 0.007 |
| Path c (Total Effect) | 0.172* | 0.029 |
| Indirect Effect (a b) | ||
| Point Estimate | 0.032 | 0.022 |
| 95% Confidence Interval | [-0.022, 0.108] | [-0.055, 0.102] |
| Significant | No | No |
| Note: . X=GenAI Literacy, M=AI-Support Score, Y=AI-Removal Score. | ||
5 Discussion
This study explored the association between GenAI literacy and students’ multimodal academic writing performance in conditions where either reactive or proactive GenAI tools supported academic writing. By examining students’ independent writing performance immediately following the withdrawal of GenAI assistance, we sought to understand whether and how GenAI literacy shapes writing performance once external scaffolding is removed and how interactions with chatbots of different styles might moderate these relationships.
5.1 GenAI Literacy and Independent Writing Performance (RQ1)
For RQ1, our findings indicated a significant positive relationship between students’ GenAI literacy and their academic writing performance at the independent writing stage immediately after GenAI assistance was removed. Specifically, participants with higher GenAI literacy demonstrated stronger performance across multiple rubric dimensions (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking, and composite scores) when writing without AI support. Importantly, these associations were robust when controlling for domain knowledge, and VLAT scores themselves showed no significant associations with writing scores. This pattern indicates that the observed effects are not attributable to differences in domain knowledge but instead reflect a unique skill set tied to GenAI literacy. This result provides empirical evidence underscoring the role of GenAI literacy as potentially influential in learners’ capabilities, particularly when performing independently without technological scaffolds. It appears that GenAI literacy is associated with superior performance in conditions where no external AI prompts or support mechanisms are immediately available. This finding is in line with prior theoretical positions suggesting that effective interactions with technology depend not only on the availability of tools but also on users’ understanding of how best to integrate and leverage them effectively (Nguyen et al., 2024; Yan et al., 2024a). Researchers argue that higher levels of literacy related to emerging digital tools lead to better competence, critical engagement with technology-generated output, and stronger independent strategic thinking (Jin et al., 2024; Kim et al., 2025). Our results reinforce this perspective in the context of GenAI-assisted multimodal writing, highlighting that students with higher GenAI literacy appear better prepared, at least immediately following GenAI assistance, to produce high-quality, independent academic writing without continual external support.
5.2 The Role of Chatbot Interaction Strategies (RQ2)
Interestingly, for RQ2, significant relationships between higher GenAI literacy and independent academic writing performance emerged only within the reactive (reactive) chatbot condition. In contrast, the proactive (actively scaffolding) chatbot condition did not demonstrate a significant correlation between GenAI literacy and subsequent independent writing performance. In addition, supplementary Mann–Whitney analyses confirmed that these results were not explained by domain knowledge, as the comparisons of high- and low-VLAT groups revealed no significant differences in writing performance across either condition or time point. This strengthens the interpretation that it was the chatbot interaction strategy, rather than differences in subject knowledge, that shaped whether GenAI literacy translated into stronger independent writing performance. This finding suggests that chatbot interaction strategies can potentially moderate the relationship between GenAI literacy and student outcomes.
One plausible explanation is that reactive chatbots, which respond only when students actively query them, require greater cognitive and strategic engagement from learners. GenAI-literate students in such conditions may have had opportunities and abilities to formulate effective prompts, critically evaluate the AI-generated assistance, and transfer the insights gained into their independent writing tasks. Conversely, proactively scaffolding chatbots continuously direct and structure learner engagement, perhaps minimising the need for students to exercise or rely upon their own GenAI literacy skills explicitly. This proactive scaffolding likely enabled both high and low GenAI literacy students to allocate their cognitive resources more effectively during task completion, thus levelling performance differences related to GenAI literacy. Students in the proactive condition benefitted from structured support, reducing the cognitive demands associated with engaging directly with GenAI tools, and enabling those with lower initial GenAI literacy to perform comparably to students with higher sophistication. Previous studies suggest scaffolding can enhance learner performance by decreasing unnecessary cognitive load, especially for those less adept with novel technologies (Nguyen et al., 2024; Sharma and Hannafin, 2007). Additionally, by providing structured guidance to learners with varying technological familiarity, this scaffolding approach may help mitigate the risk of a digital divide (Celik, 2023). However, it is also possible that the proactive condition limited opportunities for learners to actively practice and further develop their GenAI literacy skills, a consideration that warrants future investigation. Consequently, educators should balance structured support from GenAI chatbots with opportunities for independent inquiry, encouraging active and strategic engagement that leverages learners’ pre-existing technological literacy.
5.3 Predictive Role of GenAI Literacy in Multimodal Writing (RQ3)
Ordinal logistic regression (RQ3) analyses further substantiated that GenAI literacy significantly predicted students’ independent writing outcomes. Even after controlling for domain-specific content knowledge, initial AI-assisted writing scores, and chatbot interaction conditions, our analyses indicated GenAI literacy uniquely predicted stronger independent writing outcomes in visual data integration, critical thinking, and overall composite scores post-scaffolding removal. By contrast, domain-specific knowledge and chatbot conditions did not significantly predict independent performance outcomes when considered alongside GenAI literacy. These findings reinforce recent arguments positioning digital tool proficiency (especially with GenAI) as a distinct skill set separate from traditional academic knowledge bases (Jin et al., 2024; Yan, 2023; Wang, 2025). Our findings suggest that training and instruction related explicitly to GenAI literacy may effectively support students in independently performing multimodal, complex writing tasks. This unique predictive influence prompts reconsideration of GenAI literacy as a key pedagogical target, warranting purposeful curriculum inclusion rather than simple assumptions of reactive skill acquisition.
5.4 Direct Effect of GenAI Literacy (RQ4)
Our fourth research question examined whether the effect of GenAI literacy on independent writing was mediated by in-task success. The mediation pathway was not significant in either condition. This finding challenges the view that the benefits of GenAI literacy are realised mainly through improved tool-assisted performance.
The results instead point to a direct and foundational competence. In the reactive condition the indirect path was not significant, while the direct path was reliable. This suggests that GenAI literacy does not simply improve the AI-supported attempt but provides students with transferable skills they can apply independently. A possible concern is that GenAI literacy may reflect other literacies (Stolpe and Hallström, 2024), such as general writing ability, information literacy, language proficiency, or digital literacy (Yang, 2022). Three pieces of evidence in our data argue against this. First, domain knowledge relevant to the task, measured as visualisation literacy, showed no reliable association with outcomes and did not reduce the effect of GLAT when controlled. Second, the direct rather than indirect path was observed, which is inconsistent with the view that GLAT only enhances the AI-assisted attempt. Third, GLAT predicted performance only when learner agency was required in the reactive chatbot condition. This pattern is consistent with a construct that captures strategic planning, critical evaluation and metacognitive monitoring specific to working with GenAI, rather than general language or domain-specific skills. We therefore interpret GenAI literacy as a distinct, task-general competence that becomes visible when interaction design requires learners to self-initiate and regulate their engagement with GenAI (Yan et al., 2024a).
This interpretation also fits the null effects in the proactive condition. Structured scaffolding reduced the need for students to exercise these strategic capacities and equalised outcomes regardless of GenAI literacy. The influence of GenAI literacy therefore appears less about transfer from an AI-supported task and more about a latent competency that becomes important when learner agency is emphasised.
5.5 Implications for Research and Practice
The findings highlight the importance of interaction design in shaping how GenAI literacy translates into academic outcomes. The absence of any GLAT effect in the proactive condition indicates that structured scaffolding can equalise performance across learners with different levels of GenAI literacy. In contrast, when support was reactive, students’ own literacy became decisive, with higher-literacy students demonstrating stronger independent writing. This agency gap highlights two complementary levers available to educators and designers: (i) proactive scaffolding could reduce reliance on GenAI literacy during supported work, while (ii) cultivation of GenAI literacy could strengthen students’ capacity to perform independently once external support is removed.
For instructional practice, these results suggest the value of phasing GenAI-supported learning experiences. Educators might strategically integrate AI-assisted tools with fading levels of scaffolding, thereby ensuring students benefit from structured support initially but are progressively encouraged to self-initiate, monitor, and regulate their engagement. Such designs align with established scaffolding frameworks and the Zone of Proximal Development (Vygotsky and Cole, 1978; Van de Pol et al., 2010). Importantly, chatbot design choices, such as whether to respond reactively or scaffold proactively, should be made deliberately, balancing equitable support with opportunities for learner autonomy (Nguyen et al., 2024; Yan et al., 2024a; Xie et al., 2024). At the same time, the direct role of GenAI literacy points to the need for curriculum integration. Students do not uniformly acquire the strategic skills required to effectively prompt, evaluate, and refine AI-generated content (Jin et al., 2024; Nguyen et al., 2024). Embedding instruction on these literacies into academic writing curricula could enhance immediate performance while also prepare learners for transfer to other academic tasks.
Finally, there are important avenues for future research. Longitudinal studies are needed to trace how GenAI literacy develops over time and whether gains transfer across domains and task types (Wang, 2025; Kim et al., 2025). Further work should also investigate the cognitive and metacognitive processes students draw upon when engaging with AI tools (Yang et al., 2025; Cheng et al., 2024), thereby advancing theoretical models of human–AI collaboration in learning.
5.6 Limitations and Future Directions
Several limitations warrant attention when interpreting the present findings. First, the current study assessed students’ writing performance solely at two temporally adjacent points (with and immediately after the removal of AI scaffolding) within the same experimental session, limiting our ability to make broader claims about delayed or long-term improvements. Future longitudinal studies involving follow-up measures conducted at various time intervals can more definitively establish delayed effects or longer-term skill consolidation. Second, the task’s particular context, healthcare simulations involving multimodal academic writing and visual data interpretation, limits direct generalisability across disciplines or other academic writing types. Replicating or expanding this approach to diverse fields (such as STEM, social sciences, and the humanities) would clarify the robustness and generalisability of current findings. Third, although we distinguished reactive from proactive chatbot conditions, additional chatbot features (e.g., specific scaffolding strategies, transparency, adaptivity, fairness) likely shape interactions in complex ways, warranting follow-up studies that evaluate nuanced chatbot design variables. Fourth, qualitative evidence clarifying students’ cognitive processes, metacognitive reflections, and explicit interactions with GenAI tools was beyond our current analysis scope. Future research integrating qualitative approaches (think-aloud protocols, interviews, reflection statements) or detailed log-file analyses of user-AI interactions could richly explain underlying adjustments or strategic variations driven by differing GenAI literacy levels. Finally, we did not directly measure several related literacies (e.g., information/digital literacy, baseline academic writing proficiency). Although our controls (VLAT, AI-support scores) and robustness checks reduce the risk that GLAT effects reflect a proxy for other literacies, residual confounding cannot be fully ruled out. Future work should adopt multitrait–multimethod designs (e.g., combining GLAT, academic writing assessments, and digital/information literacy scales) and incorporate language proficiency screens to strengthen discriminant validity.
6 Conclusion
This paper examined how GenAI literacy shapes students’ multimodal academic writing with and without AI support. We conducted an experimental study with 79 higher education students who wrote with either reactive or proactive chatbot assistance followed by an independent task without AI support. Results showed that GenAI literacy predicted independent writing only in the reactive condition, where students had to self-initiate and regulate their engagement, but not in the proactive condition, where scaffolding equalised outcomes. Regression confirmed that this effect was robust beyond domain knowledge and prior AI-assisted performance, and mediation analysis revealed no indirect pathway, highlighting a direct influence of GenAI literacy. We introduce the concept of the agency gap to capture this pattern and argue that GenAI literacy is a distinct, task-general competence whose relevance depends on interaction design. The study contributes theoretically by clarifying when and how GenAI literacy matters, and practically by identifying two levers for equitable design: cultivating literacy to foster independence and deploying scaffolds to mitigate disparities.
Acknowledgments
Hidden for review purposes.
Data availability
Data will be made available on request.
Appendix A Generative AI Literacy Assessment Test (GLAT)
| Item | Dimension | Question | Options (Answer Bolded) |
| 1 | Know & Understand | Which of the following best describes "Generative AI"? |
A. AI that creates new content like text, images, or music by learning from existing data.
B. An AI system designed to enhance the speed and accuracy of data retrieval in search engines. C. A form of artificial intelligence that focuses on translating languages in real-time. D. AI technology used primarily for managing and organizing large databases. |
| 2 | Know & Understand | Which of the following statements best describes an LLM (Large Language Model)? |
A. It generates text by analyzing and summarizing large volumes of web content.
B. It generates text by predicting the next word based on the context of previous words. C. It generates text by translating input text into multiple languages simultaneously. D. It generates text by using pre-defined templates and filling in the blanks. |
| 3 | Know & Understand | Which of the following tasks can Generative AI perform with a high degree of accuracy? |
A. Predicting stock market trends
B. Making ethical decisions in complex scenarios C. Diagnosing rare diseases D. Generating human-like text based on prompts |
| 4 | Know & Understand | In the context of Generative AI, what is "zero-shot learning"? |
A. Training a model without any data.
B. The ability of a model to perform a task without any task-specific training. C. A method of reducing the model’s training time to zero. D. A technique for generating synthetic training data. |
| 5 | Know & Understand | Which of the following is a potential challenge when using prompt-based development for text generation? |
A. The language model can only generate binary outputs.
B. The need for extensive labeled data to train the model. C. Crafting a prompt that accurately captures the desired context and nuances. D. The requirement for complex feature engineering. |
| 6 | Know & Understand | What does the term "token" refer to in the context of a large language model (LLM)? |
A. A token is a unit of text, such as a word or a subword, that the model processes individually.
B. A token is a unique identifier assigned to each user interacting with the language model. C. A token is a security measure used to authenticate API requests to the language model. D. A token is a reward given to users for contributing valuable data to train the language model. |
| 7 | Know & Understand | Which of the following is NOT a requirement for an AI to be considered artificial general intelligence (AGI)? |
A. The ability to learn and adapt to new tasks without human intervention.
B. The capability to perform tasks across various domains with human-like proficiency. C. The ability to predict future events with perfect accuracy. D. The capacity to understand and generate natural language. |
| 8 | Know & Understand | How does RAG (Retrieval-Augmented Generation) enhance the capabilities of an LLM? |
A. By improving its grammar and syntax.
B. By providing it with real-time and relevant data. C. By increasing its computational speed. D. By enabling it to understand multiple languages. |
| 9 | Use & Apply | When using generative AI to create a marketing pitch, which of the following strategies is least likely to be effective? |
A. Supplying the AI with information about the target audience
B. Asking the AI to include unique selling points and benefits C. Requesting the AI to use persuasive language techniques D. Providing the AI with a list of competitors’ products |
| 10 | Use & Apply | After deploying a customer service chatbot, you notice that it frequently provides outdated information about company policies. What is the best course of action to address this issue? |
A. Implement a feedback loop where users can flag outdated information for review.
B. Schedule regular updates to the chatbot’s training data to include the latest company policies. C. Set up a system where complex or policy-related queries are escalated to human agents for accurate responses. D. Conduct a comprehensive audit of the chatbot’s performance metrics to identify areas for improvement. |
| 11 | Use & Apply | Suppose you have a large dataset of emails and you want to build an application to answer questions based on this dataset. Which of the following scenarios best illustrates the advantage of using RAG over prompting (i.e., without RAG)? |
A. You need to generate creative writing pieces based on the email content.
B. You want to ensure the model can answer questions even if it has never seen similar questions before. C. You need to answer questions that require specific information from different parts of the email dataset. D. You want to reduce the size of the language model to save computational resources. |
| 12 | Evaluate & Create | As a student using a Large Language Model (LLM) to gather information for an assignment, how should you approach the information it provides? |
A. The LLM’s answers are always more trustworthy than any information you will find on the internet, so you can use them without further verification.
B. The LLM’s answers are generally more trustworthy than internet sources, but you should still verify the information with other reliable sources. C. The LLM’s answers are not necessarily more trustworthy than internet sources, and you should cross-check the information with other credible references. D. The LLM’s answers are less trustworthy than internet sources because it relies on outdated information. |
| 13 | Evaluate & Create | It is unlikely for an LLM to provide an accurate summary of the latest financial market trends in real-time. Is this statement true or false? |
A. True, because the LLM’s data may be outdated due to its knowledge cutoff.
B. True, because the LLM is not good at handling numbers and structured data. C. False, because the LLM frequently updates its knowledge base. D. False, because the LLM is capable of synthesizing the latest market data automatically. |
| 14 | Evaluate & Create | A generative AI tool has provided a summary of a research paper. The summary states, "The study found that increased screen time is directly correlated with decreased attention spans in children aged 8-12." What is your next step? |
A. Accept the summary as accurate because AI tools are generally reliable.
B. Ask the AI to provide more details about the study’s methodology and results. C. Cross-check the summary with the original research paper. D. Use another AI tool to generate a summary for comparison and evaluate the consistency between both summaries |
| 15 | Evaluate & Create | While reviewing a video of a well-known public figure making controversial statements, which characteristic confirms the video was NOT generated by AI? |
A. The public figure’s voice sounds like themselves.
B. The video has a professional and polished appearance. C. The video is high-quality with smooth transitions. D. None of the above. |
| 16 | Ethics | When a generative AI system is used for screening job applications, what issue might arise concerning the quality and fairness of hiring decisions? |
A. The AI system might overlook applicants’ unique achievements and extracurricular activities.
B. The AI system could misinterpret minor formatting differences in resumes. C. The AI system might not effectively handle applications submitted in various languages. D. The AI system could reinforce existing biases found in historical hiring data. |
| 17 | Ethics | In a healthcare startup, an accurate AI model recommends treatments, but doctors don’t trust it because they can’t understand how the model arrived at its conclusions. What core issue does this scenario illustrate? |
A. The AI model uses obsolete training data.
B. The training dataset lacks sufficient diversity. C. The treatment guidelines input are incorrect. D. The AI model behaves as a black box. |
| 18 | Ethics | What are the potential copyright implications for a journalist using an AI-generated image in a commercial article? |
A. The journalist needs to check the licensing policy of the AI tool they used.
B. The AI-generated image is automatically free to use without any restrictions. C. The journalist must pay a standard licensing fee to use the AI-generated image. D. The image cannot be used in any commercial context because it is AI-generated. |
| 19 | Ethics | Should we impose restrictions on the outputs of generative AI technologies? |
A. Yes, to reduce the computational resources required for operating these technologies.
B. Yes, to prevent the dissemination of harmful or misleading content. C. No, as it would hinder technological innovation and creativity. D. No, because users should have the freedom to access all generated content. |
| 20 | Ethics | Sending personal information to cloud-based generative AI tools has little privacy concerns. |
A. True, as this information is encrypted using sophisticated algorithms during the transmission process.
B. True, as generative AI tools are black-box systems and cannot output personal information even if it is used for model training. C. False, as generative AI tools train on unencrypted data and can output private information based on their probabilistic nature. D. False, as advancements in quantum computing can easily decipher the encrypted data. |
Appendix B Mini-Visualization Literacy Assessment Test (VLAT)
| Item ID | Visualization Type | Question |
| 1 | Line Chart | What was the price of a barrel of oil in February 2020? |
| 2 | Bar Chart | What is the average internet speed in Japan? |
| 3 | Stacked Bar Chart | What is the cost of peanuts in Seoul? |
| 4 | 100% Stacked Bar Chart | Which country has the lowest proportion of Gold medals? |
| 5 | Pie Chart | What is the approximate global smartphone market share of Samsung? |
| 6 | Histogram | What distance have customers traveled in the taxi the most? |
| 7 | Scatterplot | There is a negative linear relationship between the height and the weight of the 85 males. |
| 8 | Area Chart | What was the average price of pound of coffee beans in October 2019? |
| 9 | Stacked Area Chart | What was the ratio of girls named “Isla” to girls named “Amelia” in 2012 in the UK? |
| 10 | Bubble Chart | Which city’s metro system has the largest number of stations? |
| 11 | Choropleth Map | In 2020, the unemployment rate for Washington (WA) was higher than that of Wisconsin (WI). |
| 12 | Treemap | eBay is nested in the Software category. |
Appendix C Rubric for Evaluating Academic Writing with Visual Data
| Metric | Score | Description and Example |
| Insightfulness | Level 5 (Excellent) |
- Provides highly original and profound insights.
- Provides a holistic response that addresses every key element of the analysis. - Draws thought-provoking, actionable conclusions that extend beyond obvious interpretations. Example: "Nurse 1’s disproportionate focus on Amy (50%) coupled with their central communication role and elevated stress levels (111 BPM near Bed 4) indicates an unsustainable workload distribution. Implementing a more balanced task allocation could enhance team efficiency and reduce individual stress." |
| Level 4 (Very Good) |
- Offers detailed and valid analysis with some synthesis of data.
- Provides a well-rounded response but may lack depth in one or two areas or overlook minor elements. - Conclusions are solid but may lack the depth or originality of Level 5. Example: "While focusing on Amy aligns with priorities, Nurse 1’s increased stress due to communication demands suggests that sharing responsibilities might alleviate stress and improve outcomes." |
| Level 3 (Good) |
- Provides accurate analysis based on direct observations.
- Limited synthesis of data; basic, isolated analysis with little connections. - Insights are correct but straightforward and may be somewhat obvious. Example: "Nurse 1 spent more time attending to Amy, which led to higher stress levels compared to Nurse 2." |
| Level 2 (Fair) |
- Makes surface-level observations with minimal analysis.
- Omits key elements or provides superficial explanations. - Demonstrates a limited understanding of the scenario. Example: "Nurse 1 worked on Amy for half the time and had a higher heart rate." |
| Level 1 (Poor) |
- Provides an incomplete or highly superficial response.
- Lacks any meaningful interpretation or insight. Example: "Nurse 1 was stressed with a higher heart rate while spending time on Amy." |
| Visual Data Integration | Level 5 (Excellent) |
- Fully integrates all relevant visual data types into the analysis.
- Draws meaningful connections between different data sources. - Builds a cohesive narrative that supports complex conclusions. Example: "The solo focus (50%) on Amy, Nurse 1’s central role in communication, and elevated heart rate (111 BPM at Bed 4) collectively highlight an imbalance in task distribution, emphasising the need for better teamwork." |
| Level 4 (Very Good) |
- Effectively incorporates multiple types of visual data.
- Makes connections between data types, though some integrations may be less thorough. - Uses visual data to support analysis, but may not fully exploit all available information. Example: "Nurse 1’s extensive time with Amy and higher stress levels near Bed 4 suggest that task allocation affected their stress, as shown in the charts and heart rate data." |
| Level 3 (Good) |
- Utilises at least one type of visual data effectively.
- References other data types but with limited integration. - Analysis supported by visual data but lacks depth in synthesis. Example: "According to the communication network, Nurse 1 talked more, which might explain their higher stress level." |
| Level 2 (Fair) |
- Mentions visual data superficially without meaningful integration.
- Relies on single data points without connecting them to the broader context. Example: "The chart shows Nurse 1 had a heart rate of 111 BPM near Amy." |
| Level 1 (Poor) |
- Fails to incorporate visual data into the analysis.
- May misinterpret or ignore available data. Example: "Nurse 1 was stressed, as seen in the data." |
| Organisation and Coherence | Level 5 (Excellent) |
- Presents ideas in a clear, logical order with excellent structure.
- Uses seamless transitions between points and paragraphs. - Arguments are well-developed and interconnected. Example: "By allocating her solo efforts (50%) to Amy, Nurse 1 became central in patient care. This, combined with her elevated stress levels (111 BPM near Bed 4) and primary communication role, indicates a need for task redistribution to promote balance and reduce burnout." |
| Level 4 (Very Good) |
- Maintains a clear and logical structure with minor lapses.
- Uses transitions effectively, though some areas may be less smooth. - Ideas are well-organised and mostly easy to follow. Example: "Nurse 1’s higher stress, as indicated by heart rate data, relates to her focus on Amy. Improving teamwork could help mitigate this issue." |
| Level 3 (Good) |
- Displays basic organisation but may have occasional clarity issues.
- Transitions between ideas may be abrupt or inconsistent. - Arguments are understandable but may lack depth. Example: "Nurse 1 worked more on Amy and had higher stress. Teamwork might have helped reduce her stress." |
| Level 2 (Fair) |
- Ideas are poorly organised or disjointed.
- Lacks clear transitions, making it hard to follow the argument. - Writing may be repetitive or off-topic. Example: "They worked alone a lot. Nurse 1’s stress was high. Communication is important." |
| Level 1 (Poor) |
- Lacks coherent structure; ideas appear random or jumbled.
- Writing is confusing and difficult to understand. Example: "Nurse work on Amy. Stress high. Data shows tasks." |
| Linguistic Quality | Level 5 (Excellent) |
- Writing is fluent with varied and precise vocabulary.
- Virtually free of grammatical errors. - Complex ideas are expressed clearly and elegantly. Example: "The disproportionate workload on Nurse 1, evidenced by her solo focus on Amy and central communication role, resulted in elevated stress levels. A more equitable task distribution could mitigate stress and improve patient care." |
| Level 4 (Very Good) |
- Generally clear and correct writing with minor errors.
- Vocabulary is appropriate but may lack variety. - Ideas are communicated effectively. Example: "Nurse 1’s higher stress relates to spending more time on Amy and communicating more. Sharing tasks might reduce her stress." |
| Level 3 (Good) |
- Writing is understandable but contains noticeable errors.
- Vocabulary and sentence structure are simple or repetitive. - Errors do not significantly impede understanding. Example: "Nurse 1 worked more with Amy and was more stressed than Nurse 2." |
| Level 2 (Fair) |
- Frequent grammatical errors that may hinder comprehension.
- Limited vocabulary; word choices may be inappropriate or incorrect. - Ideas are difficult to follow due to language issues. Example: "Stress Nurse 1 shows from Amy is cause by talks more and heart rate." |
| Level 1 (Poor) |
- Writing is riddled with errors; meaning is often obscured.
- Extremely limited vocabulary and flawed sentence structure. - Ideas are not effectively communicated. Example: "Nurse stress high. Work Amy. Data show." |
| Critical Thinking | Level 5 (Excellent) |
- Demonstrates deep analysis and understanding of complexities.
- Addresses possible counterarguments or alternative interpretations. - Provides well-reasoned, innovative solutions or recommendations. Example: "While prioritising Amy is essential, Nurse 1’s elevated stress (111 BPM) and communication burden indicate an imbalance. Introducing a shared workload model could distribute tasks more evenly, potentially improving team performance and reducing individual stress." |
| Level 4 (Very Good) |
- Analyses data carefully with awareness of underlying complexities.
- Offers logical conclusions and considers some implications. - Suggestions are practical and grounded in the data. Example: "Nurse 1’s stress suggests that although focusing on Amy is appropriate, collaborating more with colleagues might alleviate stress." |
| Level 3 (Good) |
- Provides basic analysis but does not delve into complexities.
- Conclusions are valid but may be obvious or lack depth. - Limited consideration of broader implications or solutions. Example: "Because Nurse 1 worked more on Amy, she was more stressed." |
| Level 2 (Fair) |
- Offers superficial analysis with little critical engagement.
- Conclusions may be unsupported or overly simplistic. Example: "Nurse 1 was stressed because she focused on Amy." |
| Level 1 (Poor) |
- Lacks critical analysis; may misinterpret data.
- Conclusions are incorrect or irrelevant. Example: "Nurse 1 stressed due to work on Amy. Nurse 2 worked less." |